### v2

# Yaohui Li2, Jun Lan1, Huijia Zhu1, Jianfu Zhang3, Weiqiang Wang1,Huaxiong Li2 1Ant Group

Abstract
Recently, video generation techniques have advanced rapidly. Given the popularity of video content on social media platforms, these models intensify concerns about the spread of fake information. Therefore, there is a growing demand for detectors capable of distinguishing between fake AI-generated videos and mitigating the potential harm caused by fake information. However, the lack of large-scale datasets from the most advanced video generators poses a barrier to the development of such detectors. To address this gap, we introduce the first AI-generated video detection dataset, GenVideo. It features the following characteristics: (1) a large volume of videos, including over one million AI-generated and real videos collected; (2) a rich diversity of generated content and methodologies, covering a broad spectrum of video categories and generation techniques. We conducted extensive studies of the dataset and proposed two evaluation methods tailored for real-world-like scenarios to assess the detectors’ performance: the cross-generator video classification task assesses the generalizability of trained detectors on generators; the degraded video classification task evaluates the robustness of detectors to handle videos that have degraded in quality during dissemination. Moreover, we introduced a plug-and-play module, named Detail Mamba (DeMamba), designed to enhance the detectors by identifying AI-generated videos through the analysis of inconsistencies in temporal and spatial dimensions. Our extensive experiments demonstrate DeMamba’s superior generalizability and robustness on GenVideo compared to existing detectors. We believe that the GenVideo dataset and the DeMamba module will significantly advance the field of AI-generated video detection. Our code and dataset will be aviliable at https://github.com/chenhaoxing/DeMamba.
1 Introduction
Advancements in generative models (Zhang et al., 2023a; Chen et al., 2023a; Li et al., 2023) have been impressive, enabling the creation of highly realistic images with less effort and expertise. As these models become capable of generating sufficiently realistic images, more researchers are exploring how to improve video creation (Blattmann et al., 2023a; Liu et al., 2023a; Wang et al., 2023a; GoogleAI, ). Currently, certain generative algorithms, such as Sora (Brooks et al., 2024) and Gen2 (Research, 2023), are capable of producing high-quality videos through the use of straightforward inputs, including text and images. While these generative algorithms can reduce manual labor and enhance creativity, they also introduce risks Barrett et al. (2023). For example, they could be utilized to misinform the public in critical domains such as politics or economics. A notable incident involved an AI-generated video of Taylor Swift that spread widely on Twitter, harming her reputation. This situation highlights the pressing need for technology that can detect these fake videos and avoid potential harm.
To assist in developing robust and highly generalizable detectors, we have created the first million-scale dataset of AI-generated videos, named GenVideo. GenVideo leverages state-of-the-art models to generate massive amounts of video, providing comprehensive training and validation for detectors of AI-generated videos. Unlike deepfake video datasets (Gu et al., 2021a; Xu et al., 2023a; Gu et al., 2022a) which focus on human face videos, GenVideo encompasses a broad spectrum of scene contents and motion variations, closely simulating the real-world authentication challenges posed by video generation models in various practical settings. GenVideo includes 1,078,838 generated videos and 1,223,511 real videos. The fake videos consist of those generated in-house and those collected from the internet, while the real videos come from the Youku-mPLUG (Xu et al., 2023b), Kinetics-400 (Kay et al., 2017a), and MSR-VTT (Xu et al., 2016a) datasets. Due to the scale of the data, we can prevent detectors from merely learning the content differences between real and fake videos, instead focusing on subtle signs that determine video authenticity. We propose two tasks that align with real-world detection challenges: (1) cross-generator video classification, where a trained detector is tasked with identifying videos from unseen generators; and (2) degraded video classification, where the detector assesses videos that have been degraded, such as those with low resolution, compression artifacts, or Gaussian blur. GenVideo can significantly advance the development of detectors aimed at identifying AI-generated videos in society.
In this paper, we evaluated state-of-the-art detection models (Qian et al., 2020; Tan et al., 2024; Gu et al., 2021a; Ni et al., 2022; Radford et al., 2021) on GenVideo. However, the generalization capabilities of these models are compromised due to the limitations of existing image detection methods, which cannot model temporal inconsistencies, and video detection methods, which struggle to efficiently model local spatial inconsistencies. In the show “Detail”  [^1] https://www.youtube.com/watch?v=NZmEY3n97P4, Kobe Bryant offered insights into basketball nuances, including Jason Tatum’s toe positioning when catching the ball, advice that significantly improved Tatum’s ability to penetrate defenses. As shown in Figure 1, generated videos often exhibit both spatial and temporal artifacts, and modeling only one aspect (either spatial or temporal) may not be sufficient to cover all types of artifacts. Building a detector with satisfactory generalization performance requires modeling the spatial-temporal local details. In this paper, we introduce a plug-and-play module called Detail Mamba (DeMamba), which leverages a structured state space model to capture spatial-temporal inconsistencies across different regions, thereby discerning the authenticity of videos. Extensive experiments on GenVideo demonstrate that DeMamba can be used as a plug-and-play addition to existing feature extractors, significantly enhancing the generalizability and robustness of models.
Our contributions are summarized as follows:
  • We introduce the first million-scale dataset for AI-generated video detection, GenVideo, which includes fake videos from various scenes, contents, and models.
  • We design two tasks to evaluate the performance of detectors: cross-generator video classification and degraded video classification.
  • We propose a plug-and-play detector, DeMamba, capable of modeling spatial-temporal inconsistencies. Extensive experimental results validate the generalizability and robustness of our DeMamba in identifying AI-generated videos.
Figure 1: Spatial and temporal artifacts in generated videos. We illustrate the spatial and temporal artifacts present in the generated videos. Artifacts: (a) errors in local appearance, (b) frequency inconsistency: average spectrum of video frames for real videos and fake videos generated, (c) temporal inconsistency.
2 Related works
2.1 Video generation methods
Video generation methods Henschel et al. (2024); Zhou et al. (2024) have become powerful tools for producing high-quality video content from textual or image prompts. Currently, video generation primarily encompasses two major tasks: Text-to-Video (T2V) and Image-to-Video (I2V). T2V involves inputting a text prompt to the model to generate videos based on textual instructions, while I2V aims to generate videos based on an input image, describing videos’ content or specific frames. Based on the types of these video generation methods, they can be separated into three categories: diffusion-based methods with Unets, diffusion-based methods with Transformers, and other methods.
Diffusion-Unet. In the realm of video synthesis Ma et al. (2024a); Zhang et al. (2024); Bar-Tal et al. (2024); Wei et al. (2023); Ho et al. (2022); Girdhar et al. (2023); Feng et al. (2023); Xu et al. (2023c); Hu et al. (2023); Ni et al. (2023); Girdhar et al. (2023) recent advancements in diffusion-unet based methods have demonstrated significant progress. Text2Video-ZeroKhachatryan et al. (2023) introduces a cost-effective approach, enriching latent codes with motion dynamics for temporal consistency. Animatediff Guo et al. (2023a) introduces a plug-and-play motion module for animating personalized text-to-image models without model-specific tuning, while Pia Zhang et al. (2023b) designs a personalized image animator for motion controllability aligned with conditioned images. Similarly, Lavie Wang et al. (2023b) leverages temporal self-attentions and rotary positional encoding, emphasizing joint image-video fine-tuning for high-quality video outputs. I2VGen-XL Wang et al. (2023c) makes preliminary attempts in video generation with visual guidance from images. VideoCrafter Chen et al. (2023b, 2024) and DynamiCrafter Xing et al. (2023) enhance video quality through synthesized images and leveraging motion priors from text-to-video models to animate open-domain images. ModelScopeT2V Wang et al. (2023d) incorporates spatial-temporal blocks to ensure consistent frame generation and smooth movement transitions. SVD Blattmann et al. (2023b) evaluates the critical multi-stage training required for successful video latent diffusion models. VideoComposer Wang et al. (2023e) creatively concatenates image embeddings with style embeddings to enhance the visual continuity of generated videos. SEINE Chen et al. (2023c) proposes a random-mask video diffusion model to push the boundaries of text-driven video synthesis.
Diffusion-Transformer. In the evolving field of video generation, diffusion-transformer based methods (Brooks et al., 2024; Ope, 2024) have garnered considerable attention due to their flexibility and efficiency in handling sequential data. Latte (Ma et al., 2024b) enhances the transformer methodology by extracting spatio-temporal tokens from input videos and modeling video distribution in the latent space using a series of Transformer blocks. Cogvideo (Hong et al., 2022) leverages a transformer-based model optimized with a multi-frame-rate hierarchical training strategy, which enhances learning efficiency and video quality. Sora (Ope, 2024; Brooks et al., 2024) adopts a DiT-based generative architecture (Li et al., 2022), for video generation, showcasing the versatility of transformer-based models in adapting to the specific demands of video synthesis.
Others. In addition to diffusion-based models, Generative Adversarial Network(GAN) Shen et al. (2023); Wang et al. (2023f) and autoregressive models (GoogleAI, ; Kondratyuk et al., 2023) are also applied for video generation. Notable contributions in this domain include(Yoo et al., 2023) and (Lei et al., 2023), who have explored the fundamentals of transformer applications in video generation. FlashVideo (Lei et al., 2023) focuses on accelerating transformer models for video generation. VideoPoet (Kondratyuk et al., 2023) utilizes a decoder-only transformer architecture to process multimodal inputs and generate creative videos. Magvit (Yu et al., 2023) employs a masked generative video transformer that quantizes videos into spatial-temporal visual tokens using a 3D tokenizer. A few works (Shen et al., 2023; Ghosh et al., 2024; Wang et al., 2023f) introduce temporal layers into GAN for video generation.
2.2 AI-Generated content detection
AI-generated visual content may raise concerns about the spread of misinformation. As a result, considerable efforts have been made to design forgery detection models and establish benchmarks in this field. In recent years, a significant amount of research has focused on detecting generated images (Guo et al., 2023b; Lorenz et al., 2023; Wu et al., 2023; Wang et al., 2023g) with help of AI-Generated Image datasets (Wang et al., 2022; Zhu et al., 2023a; Wang et al., 2023g), particularly those from unseen generative models. To date, studies in (Gu et al., 2021a; Xu et al., 2023a; Gu et al., 2022a) have addressed the detection of Deepfake videos, but there is a lack of research specifically dedicated to detecting generated videos on a wider range beyond human faces. We hope that this paper will make pioneering and insightful contributions to this research area.
Table 1: Statistics of real and generated videos in the GenVideo dataset.
Video Source Type Task Time Resolution FPS Length Training Set Testing Set Total Count
Kinetics-400 (Kay et al., 2017b) Real - 17.05 224-340 - 5-10s 260,232 - 1,213,511
Youku-mPLUG (Xu et al., 2023b) - 23.07 - - 10-120s 953,279 -
MSR-VTT (Xu et al., 2016b) Real - 16.05 - - 10-30s - 10,000 10,000
ZeroScope (zer, 2024) Fake T2V 23.07 1024$\times$576 8 3s 133,169 -
I2VGen-XL (Wang et al., 2023c) I2V 23.12 1280$\times$720 8 2s 61,975 -
SVD (Blattmann et al., 2023b) I2V 23.12 1024$\times$576 8 4s 149,026 -
VideoCrafter (Chen et al., 2024) T2V 24.01 1024$\times$576 8 2s 39,485 - 1,048,575
Pika (pik, 2022) T2V&I2V 24.02 1088$\times$640 24 3s 98,377 -
DynamiCrafter (Xing et al., 2023) I2V 24.03 1024$\times$576 8 3s 46,205 -
SD (Zhang et al., 2023b) T2V&I2V 23-24 512-1024 8 2-6s 200,720 -
SEINE (Chen et al., 2023c) I2V 24.04 1024$\times$576 8 2-4s 24,737 -
Latte (Ma et al., 2024b) T2V 24.03 512$\times$512 8 2s 149,979 -
OpenSora (Ope, 2024) T2V 24.03 512$\times$512 8 2s 177,410 -
ModelScope (Wang et al., 2023d) Fake T2V 23.03 256$\times$256 8 4s - 700 8,588
MorphStudio (Mor, 2023) T2V 23.08 1280$\times$720 8 2s - 700
MoonValley (moonvalley.ai, 2022) T2V 24.01 1024$\times$576 16 3s - 626
HotShot (Hot, 2023) T2V 23.10 672$\times$384 8 1s - 700
Show_1 (Zhang et al., 2023c) T2V 23.10 576$\times$320 8 4s - 700
Gen2 (Esser et al., 2023) I2V&T2V 23.09 896$\times$512 24 4s - 1,380
Crafter (Chen et al., 2023b) T2V 23.04 256$\times$256 8 4s - 1,400
Lavie (Wang et al., 2023a) T2V 23.09 1280$\times$2048 8 2s - 1,400
Sora (Brooks et al., 2024) T2V 24.02 - - -60s - 56
WildScrape T2V&I2V 24 512-1024 8-16 2-6s - 926
Total Count - - - - - - 2,294,594 19,588 2,314,182
3 GenVideo
3.1 Overviews of GenVideo
In response to the critical need for evaluating the generalizability of datasets and detectors (i.e., the capacity of training detectors to accurately recognize unseen videos from the open world) and the robustness of these detectors (i.e., their ability to maintain high performance against various corruptions to fake videos), we have developed the GenVideo dataset. This dataset is characterized by two main features:
  • Large scale: The GenVideo dataset is organized hierarchically, encompassing cross generators such as diffusion-based generators and transformer-based generators, and cross architectures within the same type of generator like different motion modules combined with the same T2I base model Guo et al. (2023a); Zhang et al. (2023b). This structure facilitates covering a broader range of generated content and producing fake videos on a larger scale. The training (resp., testing) set in GenVideo contains a total of $2,294,594$ (resp., $19,588$) video clips, comprising $1,213,511$ (resp., $10,000$) real videos and $1,081,083$ (resp., $8,588$) fake videos.
  • Diverse content: GenVideo includes a wide array of high-quality fake videos sourced from open-source websites, along with videos produced using both user-trained and officially provided pre-trained video generation models, including T2V and I2V models. The generated video content encompasses a diverse range of scenes, including landscapes, people, buildings, objects, and more. The duration of the videos is primarily between 2 to 6 seconds, and the aspect ratios of the video resolutions vary widely. This diverse collection ensures a comprehensive set of fake videos, significantly enriching the understanding of AI-generated video detection across numerous real-world contexts, and enhancing the generalizability and robustness of detectors.
Evaluation objectives. To avoid trivial detection caused by the same distribution from the same generator, as observed in previous AI-generated image detection datasets (Wang et al., 2020; Zhu et al., 2023b), we conduct two tasks to verify the performance of detection models: cross-generator generalization and degraded video classification. Cross-generator generalization refers to the model being trained on data generated by some generators and validated on unseen data generated by other generators, which is meant to test the model’s generalization ability. Degraded video classification, on the other hand, is used to validate the model’s robustness by testing its ability to recognize videos of different types of degradation.
3.2 Organization of GenVideo
The GenVideo dataset primarily consists of real videos and fake videos shown in Table 1. The real videos are mainly sourced from existing datasets related to video action dataset (Kay et al., 2017a) and video description dataset (Xu et al., 2023b, 2016b). The fake videos are obtained through external web scraping, internal generation pipelines based on open-source projects, and a number of existing video evaluation datasets (Liu et al., 2023b).
Considering the emergence of video generation models, which primarily focus on diffusion-based methods (Xing et al., 2023; Zhang et al., 2023b; Wang et al., 2023c; Blattmann et al., 2023b; Chen et al., 2024; Xing et al., 2023; Zhang et al., 2023b; Chen et al., 2023c) and methods based on auto-regressive models Ma et al. (2024b); Ope (2024), the training set of the GenVideo dataset predominantly comprises videos generated by these two popular types of algorithms shown in Table 1. Additionally, following (Liu et al., 2023b), we generate $98,377$ videos using the service provided by Pika website pik (2022). To balance the quantity ratio between real videos and fake videos, we sampled 260,232 and 953,279 video clips from the existing video datasets Kinetics-400 (Kay et al., 2017b) and Youku-mPLUG (Xu et al., 2023b), respectively, to form the white sample of the training set.
For the testing set, the real videos are sourced from the MSR-VTT dataset (Xu et al., 2016b), which is a large video description dataset. The fake videos are mainly sourced from two parts: the first part comes from the Evalcrafter benchmark (Liu et al., 2023b), which is used to assess the temporal smoothness, quality, and other metrics of different generation models. The second part of the data comes from external web scraping, covering generated videos from existing popular video generation methods (Zhang et al., 2024; Yang et al., 2024; Bar-Tal et al., 2024; Ma et al., 2024a; Wang et al., 2023e; Ren et al., 2024; Wang et al., 2023h; Qing et al., 2023; Ho et al., 2022; Ge et al., 2023; Guo et al., 2023c; Tian et al., 2024; Kondratyuk et al., 2023; Yoo et al., 2023; Ni et al., 2023; Zeng et al., 2023; Wei et al., 2023; Feng et al., 2023; Xu et al., 2023c; Hu et al., 2023; Yu et al., 2023), which includes diffusion-based methods (Blattmann et al., 2023a; Xing et al., 2023; Wang et al., 2023d; Zeng et al., 2023; Wei et al., 2023; Feng et al., 2023; Xu et al., 2023c; Hu et al., 2023), auto-regressive-based models (Kondratyuk et al., 2023; Ma et al., 2024b; Brooks et al., 2024), and other models (Yoo et al., 2023; Yu et al., 2023). This data encompasses most of the currently available video generation methods and advanced derivative methods of mainstream video generation techniques. Those scraped data are denoted as WildScape in Table 1.
3.3 Video collection details of GenVideo
We synthesize fake videos and gather real videos to construct the GenVideo dataset utilizing the hierarchical structure and the corresponding generators. It’s crucial to underline that the primary objective of an AI-generated video detection dataset is to achieve robust and generalizable detection capabilities, rather than solely focusing on video quality for assessment purposes. A diverse and large-scale collection ensures that the dataset encompasses a wide range of video categories, facilitating detailed evaluations of AI-generated video detection algorithms and their effectiveness across various contexts.
Fake video collection: The guiding principle for collecting fake videos is to ensure maximal diversity in content and generators. We prioritize generating additional fake videos using the most recent generators due to their superior quality. To collect diverse fake videos from different resources as training samples, we have established a video generation pipeline for text-to-video generation and image-to-video generation. This pipeline facilitates the production of videos using popular generative mechanisms, including diffusion-based models (Xing et al., 2023; Zhang et al., 2023b; Guo et al., 2023a; Wang et al., 2023c; Blattmann et al., 2023b; Chen et al., 2024; Xing et al., 2023; Zhang et al., 2023b; Chen et al., 2023c), transformer-based methods (Ma et al., 2024b; Ope, 2024) and service-based method Pika website (pik, 2022). We selected common categories such as humans, animals, and plants as foreground keywords, and typical scenes like “in the park” or “on the lawn” as background keywords. Leveraging a large language model (Le Scao et al., 2023), these foreground and background keywords are expanded into comprehensive textual prompts to guide text-to-video generation. For image-to-video generation, we employed various text-to-image models, including different versions of Stable Diffusion (SD (Rombach et al., 2022), SDXL (Podell et al., 2023)). Using the enriched textual prompts generated by the large language model, we created corresponding images, which were subsequently used as input for image-to-video models to generate the final videos.
To assemble a representative testing set, we investigated current video generators based on different model architectures (Blattmann et al., 2023a; Brooks et al., 2024; Ghosh et al., 2024) and scraped example videos from their projects, as illustrated by WildScape in Table 1. This includes prominent video generation models such as VideoPoet (Kondratyuk et al., 2023), Emu (Girdhar et al., 2023), and Sora (Brooks et al., 2024). Additionally, we collected videos generated by various condition-guided models (Wei et al., 2023; Feng et al., 2023; Xu et al., 2023c) that focus on social contexts and characters. We also included some non-mainstream generation algorithms, such as those based on latent flow diffusion models (Ni et al., 2023), masked generative video transformer (Yu et al., 2023), or autoregressive models (Gupta et al., 2023). This approach ensures coverage of both popular algorithms and those generating high-quality content, particularly around character-centric videos. We integrated existing video quality evaluation dataset Liu et al. (2023a) that include typical generation methods and demonstrate relatively high generation quality.
Real video collection: Considering that fake videos from generators are limited to specific domains determined by training datasets such as Kinetics-400 (Kay et al., 2017b) and Youku-mPLUG (Xu et al., 2023d), we sample parts of videos from those datasets as the real part of the GenVideo dataset. Specifically, we randomly sample $953,279$ videos from Youku-mPLUG (Xu et al., 2023d) and randomly slice 10-second segments from each video to form real samples.
4 DeMamba
4.1 Preliminaries
Structured State Space Sequence models (S4) (Gu et al., 2022b, 2021b; Smith et al., 2023) are grounded in continuous systems, facilitating the mapping of a one-dimension function or sequence, denoted as $x(t)\in\mathbb{R}^{L}$ to $y(t)\in\mathbb{R}^{L}$, via an intermediary hidden state $h(t)\in\mathbb{R}^{N}$. In a formal context, S4 leverage the subsequent ordinary differential equation to represent the input data:
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),\quad y(t)=\mathbf{C}h(t),$ (1)
[…]
$\mathbf{\bar{A}}={\rm exp}(\mathbf{\bigtriangleup A}),\quad\mathbf{\bar{B}}=\mathbf{\bigtriangleup A}^{-1}({\rm exp}(\mathbf{\bigtriangleup A}-\mathbf{I}))\cdot\mathbf{\bigtriangleup B},\quad h_{t}=\mathbf{\bar{A}}h_{t-1}+\mathbf{\bar{B}}x_{t},\quad y_{t}=\mathbf{C}h_{t}.$ (2)
Contrary to traditional models that primarily rely on linear time-invariant S4, Mamba (Gu and Dao, 2023) distinguishes itself by implementing a Selection mechanism computed with Scan for S4 (S6). Within the S6 framework, parameters $\mathbf{B}\in\mathbb{R}^{B\times L\times N}$, $\mathbf{C}\in\mathbb{R}^{B\times L\times N}$, and $\mathbf{\bigtriangleup}\in\mathbb{R}^{B\times L\times D}$ are inherently derived from the input $\mathbf{x}\in\mathbb{R}^{B\times L\times D}$, formulating an intrinsic structure for contextual perceptiveness and adaptive modulation of weights.
4.2 AI-generated video detection with DeMamba module
Overview. As illustrated in Figure 2, our proposed method comprises a feature encoder, a DeMamba block, and an MLP classification head. Specifically, we employ state-of-the-art vision encoders (e.g., CLIP (Radford et al., 2021) and XCLIP (Ni et al., 2022)) to encode the input video frames $\bm{X}^{v}\in\mathbb{R}^{3\times T\times H\times W}$ into a sequence of features, denoted by $\bm{F}\in\mathbb{R}^{T\times C\times H_{f}\times W_{f}}$, where $C$ symbolizes the channel dimensionality, and $H_{f}$, $W_{f}$ represent the spatial dimensions, i.e., height and width of the feature maps, respectively. Following this, the extracted features are spatially grouped, and the DeMamba module is applied to model the intra-group feature consistency. Finally, we aggregate the features from different groups to determine whether the input video is generated by AI.
DeMamba block. We first apply spatial consolidation: given the feature $\bm{F}$, we split it into $s^{2}$ zones along both the height and width dimension where each zone of $\bm{F}$ is denoted as $\bm{F}_{jk}\in\mathbb{R}^{T\times C\times(H_{f}/s)\times(W_{f}/s)}$, where $j,k=\{1,...,s\}$. In Figure 2, we adapt the 1D Mamba layer for handling spatial-temporal input by expanding its capability to a 3D scan. In the previous Mamba approaches (Gu and Dao, 2023; Zhu et al., 2024; Liu et al., 2024), a sweep-scan mechanism was utilized, which might not effectively capture the inherent contextual relationships between adjacent tokens. To address this limitation, we propose a continuous scan strategy for each segmented region, aimed at maintaining spatial continuity throughout the entire scanning phase. Suppose a zone consists of four spatial positions: (1,1), (1,2), (2,1), and (2,2), corresponding to the top-left, top-right, bottom-left, and bottom-right corners, respectively. The sweep-scan order is (1,1) -> (1,2) -> (2,1) -> (2,2), whereas in the continuous scan, the order is (1,1) -> (1,2) -> (2,2) -> (2,1). This method organizes spatial tokens based on their proximity and subsequently aligns them sequentially across successive frames. It facilitates the coherent integration of spatial and temporal dynamics, enhancing the capability of the model to capture complex spatial-temporal relationships. After modeling the spatial-temporal inconsistency of each partitioned region using DeMamba, we can obtain the feature $\bm{F}^{\prime}_{jk}\in\mathbb{R}^{T\times C\times(F_{h}F_{w}/s^{2}))}$, where $j,k=\{1,...,s\}$.
Classification head. To leverage more comprehensive features for classification, we aggregate both global and local features. Specifically, we temporally and spatially average the input features $\bm{F}$ before the DeMamba block to obtain the global feature $\bm{F}^{\rm global}\in\mathbb{R}^{C}$, and average pool the temporal and spatial features $\bm{F}^{\prime}_{jk}$ after the DeMamba processing into pooled features $\bm{F}^{\rm pool}_{jk}\in\mathbb{R}^{C}$. Then we concatenate the local features with the global features and apply a simple MLP for classification:
$y_{\rm pred}={\rm Sigmoid}({\rm MLP}([\bm{F}^{\rm global};\bm{F}^{\rm pool}_{11};....;\bm{F}^{\rm pool}_{ss}])).$ (3)
Finally, we use binary cross-entropy loss to train our model to classify real/fake videos.
Figure 2: The overall framework of our Detail Mamba (DeMamba).
5 Experiments
5.1 Implementation details
Datasets. To comprehensively analyze the performance of various detectors, we divided the dataset into two distinct parts: the basic training set $D_{\rm train}$ and the out-of-domain test set $D_{\rm v-ood}$. $D_{\rm train}$ and $D_{\rm v-ood}$ contain fake videos created by different generative methods and different real videos. $D_{\rm train}$ and $D_{\rm v-ood}$ contain fake videos produced by different generative methods and real videos from different sources. $D_{\rm train}$ includes 1,213,511 real videos and 1,048,575 generated videos produced using 10 baseline generative methods. $D_{\rm v-ood}$ contains 10,000 real videos and 8,588 generated videos created with 10 generative methods. For detailed information about the data, please refer to Section 3 and Section 3.3.
Table 2: Training parameter settings.
Task Model Batchsize LR Samples Epochs Optimizer Scheduler Loss
per epoch
F3Net 1024 1e-5 $1024\times 2000$ 30 AdamW [20, 25], lr * 0.1 BCE
NPR 1024 1e-5 $1024\times 2000$ 30 [20, 25], lr * 0.1
STIL 128 1e-5 $128\times 4000$ 30 [20, 25], lr * 0.1
VideoMAE-B 128 1e-5 $128\times 4000$ 30 [20, 25], lr * 0.1
many- CLIP-B-LP 1024 1e-6 $1024\times 2000$ 30 [20, 25], lr * 0.1
to- CLIP-B-FT 1024 1e-6 $1024\times 2000$ 10 -
many XCLIP-B-LP 128 1e-6 $128\times 4000$ 30 [20, 25], lr * 0.1
generalization XCLIP-B-FT 128 1e-6 $128\times 4000$ 10 -
DeMamba-CLIP-LP 128 1e-6 $1024\times 2000$ 30 [20, 25], lr * 0.1
DeMamba-CLIP-FT 128 1e-6 $1024\times 2000$ 10 -
DeMamba-XCLIP-LP 128 1e-6 $128\times 4000$ 30 [20, 25], lr * 0.1
DeMamba-XCLIP-FT 128 1e-6 $128\times 4000$ 10 -
one- STIL 128 1e-5 $128\times 400$ 10 AdamW - BCE
to- NPR 1024 1e-5 $1024\times 2000$ 10 -
many XCLIP-B-FT 1024 1e-6 $128\times 400$ 10 -
generalization DeMamba-XCLIP-FT 128 1e-6 $128\times 400$ 10 -
Table 3: Comparisons to the SOTAs in mean accuracy (ACC) and average precision (AP) on the many-to-many generalization task.
Model Detection Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real Avg.
level Studio Valley Scope Scrape
F3Net Image ACC 83.93 99.71 98.62 77.57 57.00 36.57 99.52 99.71 89.43 76.78 99.14 83.45
AP 68.27 99.89 99.67 89.35 85.24 63.17 99.58 99.89 93.80 88.41 - 88.73
NPR Image ACC 91.07 99.57 99.49 24.29 89.64 57.71 97.12 99.86 94.29 87.80 97.46 85.30
AP 67.17 99.14 99.20 22.76 93.91 61.76 96.33 99.72 94.15 90.40 - 82.45
STIL Video ACC 78.57 98.14 98.04 76.00 61.79 53.29 99.36 97.36 94.57 65.01 98.72 83.71
AP 57.21 99.08 99.32 86.19 82.24 70.43 99.25 98.96 97.18 81.32 - 87.12
VideoMAE-B Video ACC 67.86 96.00 98.41 96.14 77.14 80.43 97.44 96.93 96.29 68.36 99.71 88.61
AP 66.49 98.85 99.77 99.27 96.55 95.31 99.49 99.69 99.27 90.74 - 94.54
CLIP-B-PT Image ACC 85.71 82.43 90.36 71.00 79.29 75.43 89.62 86.29 82.14 75.16 57.22 79.67
AP 6.78 43.56 70.88 29.97 52.97 35.36 55.52 66.03 44.23 42.99 - 44.83
DeMamba Video ACC 58.93 96.43 93.12 68.00 69.36 69.00 89.14 91.86 96.14 56.59 98.06 $80.60_{{\color[rgb]{0.453125,0,0}+0.93}}$
-CLIP-PT AP 25.87 95.14 96.23 73.43 83.31 75.49 90.17 95.06 95.05 69.95 - $79.97_{{\color[rgb]{0.453125,0,0}+35.14}}$
CLIP-B-FT Image ACC 94.64 99.86 91.38 77.29 88.14 86.00 99.68 99.79 84.29 84.67 97.38 91.19
AP 80.67 99.67 95.24 82.20 93.48 88.62 99.55 99.79 86.93 89.08 - 91.52
DeMamba Video ACC 95.71 100.00 98.70 69.14 92.43 93.29 100.00 100.00 83.57 82.94 99.44 $92.29_{{\color[rgb]{0.453125,0,0}+1.10}}$
CLIP-B-FT AP 85.50 100.00 99.59 76.15 96.78 96.99 99.97 100.00 89.80 89.72 - $93.45_{{\color[rgb]{0.453125,0,0}+1.93}}$
XCLIP-B-PT Video ACC 81.34 82.15 83.35 80.98 81.82 81.55 82.14 82.98 81.93 81.10 81.37 81.88
AP 16.39 72.16 87.77 39.86 65.57 54.26 75.23 84.80 61.60 55.28 - 61.29
DeMamba Video ACC 66.07 95.86 94.64 77.86 75.36 80.29 90.89 92.50 96.00 66.41 95.12 $84.64_{{\color[rgb]{0.453125,0,0}+2.76}}$
-XCLIP-PT AP 18.26 93.50 94.72 69.94 78.08 71.50 83.95 92.23 93.54 68.10 - $76.38_{{\color[rgb]{0.453125,0,0}+15.09}}$
XCLIP-B-FT Video ACC 82.14 99.57 93.62 61.29 79.36 69.71 97.92 99.79 77.14 83.59 98.14 85.66
AP 64.42 99.73 96.78 70.98 90.35 77.28 97.34 99.84 82.01 88.97 - 86.77
DeMamba Video ACC 98.21 100.00 99.86 65.43 94.86 98.86 100.00 100.00 92.86 89.09 99.42 $\textbf{94.42}_{{\color[rgb]{0.453125,0,0}+8.76}}$
-XCLIP-FT AP 93.32 100.00 99.97 85.55 98.97 99.60 99.98 100.00 97.77 95.75 - $\textbf{97.10}_{{\color[rgb]{0.453125,0,0}+10.43}}$
Table 4: Comparisons to the SOTAs in mean accuracy (ACC) and average precision (AP) on the one-to-many generalization task.
Training Model Detection Metric Testing subset Avg.
Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real
subset level Studio Valley Scope Scrape
Pika NPR Image ACC 55.36 77.57 71.88 4.86 7.21 4.29 86.26 60.29 71.43 31.53 99.52 51.83
AP 45.74 91.55 92.71 21.80 44.32 22.74 95.04 90.03 84.91 60.88 - 64.97
STIL Image ACC 75.00 79.43 94.49 57.86 53.14 64.14 97.12 85.29 69.43 62.42 92.43 75.52
AP 22.35 71.62 93.19 40.61 53.24 47.73 94.94 85.82 58.99 61.91 - 63.04
XCLIP-B-FT Video ACC 67.86 91.29 96.23 12.00 22.36 9.14 99.84 83.43 75.57 51.84 99.64 64.47
AP 71.08 97.53 99.44 44.68 72.69 38.37 99.96 97.32 88.00 74.00 - 78.31
DeMamba Video ACC 92.86 97.29 98.48 38.29 53.50 41.43 99.84 94.07 77.29 64.15 98.65 $\textbf{77.80}_{{\color[rgb]{0.453125,0,0}+13.33}}$
-XCLIP-FT AP 77.75 98.42 99.16 52.97 76.72 56.24 99.80 97.91 82.83 74.81 - $\textbf{81.66}_{{\color[rgb]{0.453125,0,0}+3.35}}$
SEINE NPR Image ACC 46.43 78.57 63.70 21.86 7.00 3.29 92.97 89.29 33.86 24.84 99.70 51.05
AP 36.30 92.63 85.02 52.68 25.69 11.05 97.80 97.78 64.64 47.48 - 61.11
STIL Video ACC 71.43 80.43 88.48 67.71 54.57 55.71 93.93 89.57 72.00 50.11 92.27 74.20
AP 23.89 71.01 88.18 52.17 54.49 41.23 84.73 87.38 58.72 46.51 - 60.83
XCLIP-B-FT Video ACC 85.71 95.43 76.23 65.86 35.93 37.00 99.68 99.00 75.57 49.78 99.80 74.54
AP 85.89 97.97 94.40 92.81 81.68 77.68 98.48 98.91 92.27 67.91 - 88.80
DeMamba Video ACC 94.64 98.43 92.17 82.43 52.29 54.00 99.52 99.14 79.29 57.88 98.99 $\textbf{82.61}_{{\color[rgb]{0.453125,0,0}+8.07}}$
-XCLIP-FT AP 83.74 99.01 97.66 90.82 84.11 73.30 99.72 99.73 89.72 76.45 - $\textbf{89.43}_{{\color[rgb]{0.453125,0,0}+0.63}}$
OpenSora NPR Image ACC 55.36 76.29 55.51 58.57 76.50 22.43 74.92 83.07 29.86 60.37 95.95 62.62
AP 25.65 75.24 65.02 55.12 82.42 20.75 72.65 86.84 28.13 64.50 - 57.63
STIL Video ACC 32.14 45.43 56.45 35.14 45.07 34.57 57.83 63.14 19.86 43.95 98.13 48.33
AP 6.94 55.62 75.99 43.55 68.06 44.01 63.84 80.59 29.39 57.58 - 52.56
XCLIP-B-FT Video ACC 67.86 75.86 67.46 70.86 73.14 43.57 79.87 86.29 33.43 63.17 98.10 69.06
AP 48.28 81.39 81.84 77.38 86.08 51.87 83.41 93.18 39.27 72.74 - 71.54
DeMamba Video ACC 55.36 87.43 81.30 73.14 85.21 73.14 89.62 90.07 44.86 58.10 97.30 $\textbf{75.95}_{{\color[rgb]{0.453125,0,0}+6.89}}$
-XCLIP-FT AP 25.89 86.63 87.27 74.38 91.12 76.01 86.41 93.83 48.74 67.92 - $\textbf{73.82}_{{\color[rgb]{0.453125,0,0}+2.28}}$
Table 5: Robustness evaluation of different detectors on many-to-many generalization task.
Model Detection Metric Original Compression Transformation Watemarks Gaussian Color
level ${\rm CRF}=28$ JEPG Flip Crop Text Image noise transform
F3Net Image ACC 83.45 $77.55_{{\color[rgb]{0,0.453125,0}-5.90}}$ $82.04_{{\color[rgb]{0,0.453125,0}-1.41}}$ $81.45_{{\color[rgb]{0,0.453125,0}-2.00}}$ $69.71_{{\color[rgb]{0,0.453125,0}-13.74}}$ $76.90_{{\color[rgb]{0,0.453125,0}-7.55}}$ $78.67_{{\color[rgb]{0,0.453125,0}-4.78}}$ $82.94_{{\color[rgb]{0,0.453125,0}-0.51}}$ $82.93_{{\color[rgb]{0,0.453125,0}-0.52}}$
AP 88.73 $87.39_{{\color[rgb]{0,0.453125,0}-1.34}}$ $\textbf{81.02}_{{\color[rgb]{0,0.453125,0}-7.71}}$ $87.70_{{\color[rgb]{0,0.453125,0}-1.73}}$ $67.05_{{\color[rgb]{0,0.453125,0}-21.68}}$ $87.01_{{\color[rgb]{0,0.453125,0}-1.72}}$ $87.93_{{\color[rgb]{0,0.453125,0}-0.70}}$ $87.17_{{\color[rgb]{0,0.453125,0}-1.56}}$ $87.17_{{\color[rgb]{0,0.453125,0}-1.56}}$
NPR Image ACC 85.30 $80.83_{{\color[rgb]{0,0.453125,0}-4.47}}$ $73.92_{{\color[rgb]{0,0.453125,0}-11.38}}$ $83.82_{{\color[rgb]{0,0.453125,0}-1.48}}$ $71.69_{{\color[rgb]{0,0.453125,0}-13.61}}$ $83.26_{{\color[rgb]{0,0.453125,0}-2.04}}$ $82.43_{{\color[rgb]{0,0.453125,0}-2.87}}$ $84.94_{{\color[rgb]{0,0.453125,0}-0.36}}$ $85.03_{{\color[rgb]{0,0.453125,0}-0.47}}$
AP 82.45 $76.12_{{\color[rgb]{0,0.453125,0}-6.33}}$ $53.72_{{\color[rgb]{0,0.453125,0}-28.73}}$ $79.05_{{\color[rgb]{0,0.453125,0}-3.40}}$ $36.36_{{\color[rgb]{0,0.453125,0}-46.09}}$ $67.16_{{\color[rgb]{0,0.453125,0}-15.28}}$ $77.02_{{\color[rgb]{0,0.453125,0}-5.43}}$ $57.55_{{\color[rgb]{0,0.453125,0}-24.90}}$ $76.81_{{\color[rgb]{0,0.453125,0}-5.64}}$
STIL Video ACC 83.71 $78.14_{{\color[rgb]{0,0.453125,0}-5.57}}$ $73.79_{{\color[rgb]{0,0.453125,0}-9.92}}$ $81.97_{{\color[rgb]{0,0.453125,0}-1.74}}$ $77.58_{{\color[rgb]{0,0.453125,0}-6.13}}$ $77.64_{{\color[rgb]{0,0.453125,0}-6.07}}$ $74.16_{{\color[rgb]{0,0.453125,0}-9.55}}$ $83.44_{{\color[rgb]{0,0.453125,0}-0.27}}$ $83.47_{{\color[rgb]{0,0.453125,0}-0.24}}$
AP 87.12 $85.04_{{\color[rgb]{0,0.453125,0}-2.08}}$ $59.38_{{\color[rgb]{0,0.453125,0}-27.74}}$ $85.21_{{\color[rgb]{0,0.453125,0}-1.91}}$ $76.38_{{\color[rgb]{0,0.453125,0}-10.74}}$ $82.59_{{\color[rgb]{0,0.453125,0}-4.53}}$ $83.09_{{\color[rgb]{0,0.453125,0}-4.03}}$ $74.54_{{\color[rgb]{0,0.453125,0}-12.58}}$ $86.65_{{\color[rgb]{0,0.453125,0}-0.47}}$
CLIP-B-FT Image ACC 91.19 $82.92_{{\color[rgb]{0,0.453125,0}-8.27}}$ $76.78_{{\color[rgb]{0,0.453125,0}-12.41}}$ $89.23_{{\color[rgb]{0,0.453125,0}-12.41}}$ $66.91_{{\color[rgb]{0,0.453125,0}-24.28}}$ $79.65_{{\color[rgb]{0,0.453125,0}-11.54}}$ $78.58_{{\color[rgb]{0,0.453125,0}-12.61}}$ $88.50_{{\color[rgb]{0,0.453125,0}-2.69}}$ $88.81_{{\color[rgb]{0,0.453125,0}-2.38}}$
AP 91.52 $90.42_{{\color[rgb]{0,0.453125,0}-1.10}}$ $85.33_{{\color[rgb]{0,0.453125,0}-12.41}}$ $91.05_{{\color[rgb]{0,0.453125,0}-0.14}}$ $76.35_{{\color[rgb]{0,0.453125,0}-15.17}}$ $88.94_{{\color[rgb]{0,0.453125,0}-2.58}}$ $91.05_{{\color[rgb]{0,0.453125,0}-0.47}}$ $84.55_{{\color[rgb]{0,0.453125,0}-6.97}}$ $91.05_{{\color[rgb]{0,0.453125,0}-0.50}}$
XCLIP-B-FT Video ACC 85.66 $84.17_{{\color[rgb]{0,0.453125,0}-1.49}}$ $53.96_{{\color[rgb]{0,0.453125,0}-31.70}}$ $84.48_{{\color[rgb]{0,0.453125,0}-1.18}}$ $56.24_{{\color[rgb]{0,0.453125,0}-30.42}}$ $85.06_{{\color[rgb]{0,0.453125,0}-1.60}}$ $82.84_{{\color[rgb]{0,0.453125,0}-2.82}}$ $82.94_{{\color[rgb]{0,0.453125,0}-2.72}}$ $85.41_{{\color[rgb]{0,0.453125,0}-0.25}}$
AP 86.77 $86.53_{{\color[rgb]{0,0.453125,0}-0.24}}$ $54.11_{{\color[rgb]{0,0.453125,0}-32.66}}$ $86.76_{{\color[rgb]{0,0.453125,0}-0.01}}$ $46.10_{{\color[rgb]{0,0.453125,0}-40.67}}$ $86.71_{{\color[rgb]{0,0.453125,0}-0.06}}$ $86.74_{{\color[rgb]{0,0.453125,0}-0.03}}$ $68.05_{{\color[rgb]{0,0.453125,0}-14.89}}$ $86.01_{{\color[rgb]{0,0.453125,0}-0.76}}$
DeMamba Video ACC 92.29 $85.63_{{\color[rgb]{0,0.453125,0}-6.66}}$ $83.49_{{\color[rgb]{0,0.453125,0}-8.80}}$ $90.59_{{\color[rgb]{0,0.453125,0}-1.70}}$ $62.25_{{\color[rgb]{0,0.453125,0}-30.04}}$ $83.34_{{\color[rgb]{0,0.453125,0}-8.95}}$ $87.14_{{\color[rgb]{0,0.453125,0}-5.15}}$ $90.63_{{\color[rgb]{0,0.453125,0}-1.66}}$ $91.50_{{\color[rgb]{0,0.453125,0}-0.79}}$
-CLIP-FT AP 93.45 $90.93_{{\color[rgb]{0,0.453125,0}-2.52}}$ $74.03_{{\color[rgb]{0,0.453125,0}-19.42}}$ $93.44_{{\color[rgb]{0,0.453125,0}-0.01}}$ $61.12_{{\color[rgb]{0,0.453125,0}-32.33}}$ $90.08_{{\color[rgb]{0,0.453125,0}-3.37}}$ $93.35_{{\color[rgb]{0,0.453125,0}-0.10}}$ $82.99_{{\color[rgb]{0,0.453125,0}-10.46}}$ $93.39_{{\color[rgb]{0,0.453125,0}-0.06}}$
DeMamba Video ACC 94.42 $\textbf{90.52}_{{\color[rgb]{0,0.453125,0}-3.90}}$ $\textbf{93.26}_{{\color[rgb]{0,0.453125,0}-1.16}}$ $\textbf{94.26}_{{\color[rgb]{0,0.453125,0}-0.16}}$ $\textbf{72.98}_{{\color[rgb]{0,0.453125,0}-22.44}}$ $\textbf{89.59}_{{\color[rgb]{0,0.453125,0}-4.83}}$ $\textbf{91.06}_{{\color[rgb]{0,0.453125,0}-3.36}}$ $\textbf{94.36}_{{\color[rgb]{0,0.453125,0}-0.06}}$ $\textbf{93.82}_{{\color[rgb]{0,0.453125,0}-0.60}}$
-XCLIP-FT AP 97.10 $\textbf{96.00}_{{\color[rgb]{0,0.453125,0}-1.10}}$ $76.29_{{\color[rgb]{0,0.453125,0}-20.81}}$ $\textbf{96.76}_{{\color[rgb]{0,0.453125,0}-0.34}}$ $\textbf{67.18}_{{\color[rgb]{0,0.453125,0}-29.92}}$ $\textbf{94.30}_{{\color[rgb]{0,0.453125,0}-2.80}}$ $\textbf{97.07}_{{\color[rgb]{0,0.453125,0}-0.03}}$ $\textbf{85.05}_{{\color[rgb]{0,0.453125,0}-12.05}}$ $\textbf{97.05}_{{\color[rgb]{0,0.453125,0}-0.05}}$
Table 6: Ablation study of DeMamba.
Model Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real Avg.
Studio Valley Scope Scrape
w/o Demamba ACC 82.14 99.57 93.62 61.29 79.36 69.71 97.92 99.79 77.14 83.59 98.14 $85.66_{{\color[rgb]{0,0.453125,0}-8.76}}$
Block AP 64.42 99.73 96.78 70.98 90.35 77.28 97.34 99.84 82.01 88.97 - $86.77_{{\color[rgb]{0,0.453125,0}-10.43}}$
w/o global ACC 97.33 99.85 99.45 67.66 93.55 97.88 100.00 100.00 91.34 86.54 98.55 $93.83_{{\color[rgb]{0,0.453125,0}-0.59}}$
feat AP 87.32 99.35 99.22 82.44 97.99 99.42 99.42 99.85 97.32 94.98 - $95.73_{{\color[rgb]{0,0.453125,0}-1.37}}$
DeMamba ACC 98.21 100.00 99.86 65.43 94.86 98.86 100.00 100.00 92.86 89.09 99.42 94.42
-XCLIP-FT AP 93.32 100.00 99.97 85.55 98.97 99.60 99.98 100.00 97.77 95.75 - 97.10
Table 7: Influence of different zone sizes in DeMamba.
Model Zone Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real Avg.
size Studio Valley Scope Scrape
$1\times 1$ ACC 100.00 99.86 84.78 95.57 81.43 82.29 100.00 99.79 78.29 81.10 97.64 90.98
AP 93.80 99.64 91.78 96.03 90.04 86.13 99.92 99.81 82.12 86.96 - 92.62
$2\times 2$ ACC 95.71 100.00 98.70 69.14 92.43 93.29 100.00 100.00 83.57 82.94 99.44 92.29
DeMamba AP 85.50 100.00 99.59 76.15 96.78 96.99 99.97 100.00 89.80 89.72 - 93.45
-CLIP-FT $7\times 7$ ACC 98.21 100.00 98.41 65.00 94.00 95.43 100.00 100.00 80.86 84.34 98.19 91.62
AP 87.28 99.99 99.06 73.51 97.47 97.00 99.76 99.97 85.75 89.95 - 92.97
$14\times 14$ ACC 100.00 99.86 93.41 88.57 81.71 70.14 100.00 99.93 76.43 80.02 98.70 89.89
AP 94.94 99.42 97.03 93.56 94.02 84.56 99.12 99.22 77.60 89.46 - 92.89
$1\times 1$ ACC 91.07 100.00 99.35 73.71 89.50 96.29 100.00 100.00 94.71 81.21 99.37 93.20
AP 86.65 99.99 99.83 90.64 97.85 98.81 99.91 99.99 98.43 92.18 - 96.43
$2\times 2$ ACC 98.21 100.00 99.86 65.43 94.86 98.86 100.00 100.00 92.86 89.09 99.42 94.42
DeMamba AP 93.32 100.00 99.97 85.55 98.97 99.60 99.98 100.00 97.77 95.75 - 97.10
-XCLIP-FT $7\times 7$ ACC 96.43 100.00 99.57 56.00 88.71 96.00 100.00 100.00 94.43 79.48 99.64 91.84
AP 96.02 100.00 99.97 82.85 99.69 99.52 99.98 100.00 86.51 97.45 - 96.20
$14\times 14$ ACC 76.79 100.00 99.06 22.86 84.71 95.14 99.52 100.00 87.71 72.25 99.87 85.26
AP 80.43 100.00 99.76 65.69 98.49 99.31 99.79 99.91 96.87 88.31 - 92.86
Table 8: Influence of scanning order in DeMamba.
Model Scan Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real Avg.
Rrder Studio Valley Scope Scrape
sweep ACC 95.21 100.00 92.55 61.77 92.69 96.33 99.87 99.32 81.53 73.11 98.79 $90.11_{{\color[rgb]{0,0.453125,0}-3.31}}$
DeMamba AP 93.32 99.95 97.32 85.55 95.88 97.54 99.66 97.54 87.13 80.82 - $93.47_{{\color[rgb]{0,0.453125,0}-3.63}}$
-XCLIP-FT continuous ACC 98.21 100.00 99.86 65.43 94.86 98.86 100.00 100.00 92.86 89.09 99.42 94.42
AP 93.32 100.00 99.97 85.55 98.97 99.60 99.98 100.00 97.77 95.75 - 97.10
Evaluation metrics. Consistent with the methodologies employed in prior studies, our evaluation framework primarily focuses on reporting accuracy (ACC) and average precision (AP) to assess the effectiveness of the detectors. The accuracy calculation is based on a threshold value of 0.5. For image-based detection techniques, we consolidate frame-level predictions to derive the corresponding video-level predictions, ensuring a coherent analysis across different media formats. It is noteworthy that when evaluating the performance on a dataset generated by a specific synthesis method, we calculate the ACC of that synthesis method based on the dataset itself. Additionally, in the process of computing the AP, we take into account real videos to achieve a more comprehensive assessment.
Baselines. 1) CLIP (ICML’21) (Radford et al., 2021) is an innovative model that connects images with textual descriptions, enabling it to perform exceptionally well across various vision and language tasks. 2) F3Net (ECCV’20) (Qian et al., 2020) addresses the detection of sophisticated face manipulations using a combination of frequency-aware components and loc al statistics. Its dual-stream approach effectively identifies forgery patterns. 3) VideoMAE (NeurIPS’22) (Tong et al., 2022) adapts the Masked Autoencoder (He et al., 2022) concept for video, enhancing feature extraction and improving performance on downstream tasks through self-supervised training. 4) XCLIP (ECCV’22) (Ni et al., 2022) adapts existing image-language models for video recognition, adding a new cross-frame attention mechanism to improve the exchange of temporal information and tailor prompts specifically for videos. 5) STIL (MM’21) (Gu et al., 2021a) proposes a novel spatial-temporal inconsistency learning method, which captures the spatial-temporal inconsistencies in forged videos by introducing a spatial inconsistency module, a temporal inconsistency module, and an information supplementation module. 6) NPR (CVPR’24) (Tan et al., 2024) is a concept introduced to understand and describe the complex, local interdependencies between image pixels that arise from upsampling processes, particularly evident in synthetic images created by generative models like GANs and diffusion models, which are used to identify and analyze the structural artifacts resulting from these operations.
Training settings. As shown in Table 2, we present the different model training parameter settings for many-to-many and one-to-many generalization tasks. All of our experiments were conducted on a system equipped with 8 Tesla A-100 80G GPUs and an Intel(R) Xeon(R) Platinum 8369B CPU @ 2.90GHz.
5.2 Task1: cross generator generalization
Due to the rapid iteration of generation methods, we propose a cross-dataset generalization task to test the generalization performance of detectors. Specifically, it consists of two types of generalization tasks: 1) the many-to-many generalization task, and 2) the one-to-many generalization task.
Many-to-many generalization task. This task involves training on 10 baseline categories and then testing on each subset and the average detection performance on $D_{\rm v-ood}$. As shown in Table 3, video models achieve better recognition accuracy compared to image models because video models can model temporal sequences. Moreover, our DeMamba model can be effectively integrated into existing models, achieving significant improvements. For instance, by incorporating the DeMamba module into the XCLIP, the DeMamba-XCLIP-FT attains an average accuracy of 94.42% and an average AP of 97.10%, which represents a 10.22% increase in accuracy and 12.02% increase in AP compared to the original XCLIP. Note that PT (Partially Tuning) indicates that the backbone is frozen, with only the other parts being fine-tuned, while FT (Full Training) tuning the entire model.
One-to-many generalization task. Following AI-generated image detection setting (Tan et al., 2024; Corvi et al., 2023; Wang et al., 2023g), we also perform a one-to-many generalization task. Unlike the many-to-many generalization task, the one-to-many generalization task involves training on one baseline category and then testing on each subset and the average detection performance on $D_{\rm v-ood}$. As shown in Table 4, our DeMamba-XCLIP-FT achieves better generalization performance in three one-to-many generalization tasks due to the learning of spatial-temporal inconsistency in DeMamba.
5.3 Task2: degraded video classification
In practical detection scenarios, the robustness of the detector to perturbations is also of paramount importance. In this regard, we investigated the impact of perturbations on the detector on 8 different types: H.264 compression, JPEG compression, FLIP, Crop, text watermark, image watermark, Gaussian noise, and color transform. More specific details about the perturbations can be found in Section B.2. Table 5 shows the performance of the models trained in the many-to-many task under the influence of these perturbations. We can see that in the case of degraded data, DeMamba-XCLIP-FT still achieves the best performance, indicating that our model has good robustness in the face of degraded data.
5.4 Ablation study
Ablation testing. We conducted ablation experiments to validate the effectiveness of DeMamba. As shown in Table 6, DeMamba effectively enhances the generalization performance of the model. Additionally, when using fused features, the model achieves its best performance.
Influence of different zone sizes. We investigated the impact of zone size in dividing zones in DeMamba on modeling temporal inconsistency. As shown in Table 7, the best performance is observed when the zone size is 2. Smaller zones enable the model to concentrate more on local details, leading to superior modeling performance. However, excessively small zones may result in the loss of spatial contextual information. Therefore, selecting an appropriate zone size is crucial.
Influence of scanning orders. As shown in Table 8, the continuous scan proposed in this paper effectively enhances performance compared to the traditional scanning method.
6 Broader impacts
Our research focuses on utilizing machine learning to detect generated videos. We have introduced the first million-scale AI-generated video detection dataset and developed the DeMamba model. These efforts are crucial for protecting digital content and preventing the spread of misinformation. However, there is a potential for these tools to be misused, leading to competition between video generation and detection technologies. We aim to advocate for the ethical use of technology and promote creative research into tools that verify media authenticity. We believe this will help protect the public from the harm of misinformation, enhance the clarity and authenticity of information dissemination, and ensure the protection of personal privacy.
7 Conclusion and limitation
This paper introduces GenVideo, a dataset specifically designed for detecting fake videos generated by generative models. GenVideo is characterized by its large-scale nature, as well as the rich diversity of generated content and methods. We propose two tasks that mimic real-world scenarios, namely the cross-generator video classification task and the degraded video classification task, to evaluate the detection performance of existing detectors on GenVideo. Additionally, we introduce a plug-and-play effective detection model called Detail Mamba (DeMamba), which distinguishes AI-generated videos by analyzing inconsistencies in the spatial-temporal dimensions. This model has demonstrated its strong generalization and robustness across multiple tasks. We hope that this research will inspire the creation and improvement of other detection technologies, providing new avenues for the development of authentic and reliable AI-generated content applications.
The main limitation of this article lies in the suboptimal training efficiency of the proposed DeMamba, a common issue with the Mamba model. Consequently, we encourage the community to design more lightweight and generalized detection models to facilitate the regulation of AI-generated content.
Appendix
Appendix A Model details
We provide detailed information about the methods used in this paper, as shown in Table 9. Our model only requires a small addition of parameters on the XCLIP-B model to achieve significant performance improvements.
Table 9: Model details and its performance on many-to-many generalization task.
Model Detection level Input size Param (M) FLOPs(G) ACC AP
CLIP-B Image $8\times 224\times 224$ 151.46 149.34 91.19 91.52
NPR Image $8\times 224\times 224$ 1.44 14.08 85.30 82.45
F3Net Image $8\times 299\times 299$ 48.31 145.04 83.45 88.73
VideoMAE-B Video $16\times 224\times 224$ 86.54 147.52 88.61 94.54
STIL Video $8\times 224\times 224$ 22.69 38.06 83.71 87.12
XCLIP-B Video $8\times 224\times 224$ 121.26 141.28 85.66 86.77
DeMamba-XCLIP Video $8\times 224\times 224$ 125.37 147.61 94.42 97.10
Appendix B Experiment details
B.1 Implementation details
Data pre-processing. For each video, we uniformly sample frames for alignment. For videos longer than 3 seconds, we sample 2 frames every second. For videos shorter than 3 seconds, our sampling frequency is $\frac{8}{\rm length}$ seconds. The pseudo-code for Pytorch-like is as follows:
1 video_length = get_video_length(video_path)
2 os.makedirs(os.path.dirname(image_path), exist_ok=True)
3 if video_length >= 3 :
4 inter_val = 2
5 os.system(f"cd {image_path} | ffmpeg -loglevel quiet -i {video_path}
6 -r {inter_val} {image_path}%d.jpg")
7 else:
8 inter_val = math.ceil(8 / video_length)
9 os.system(f"cd {image_path} | ffmpeg -loglevel quiet -i {video_path}
10 -r {inter_val} {image_path}%d.jpg")
Dataset augmentation. During training and testing, we randomly select 8 or 16 consecutive frames from the video after frame sampling, and resize each frame to $224\times 224$. To enhance generalizability of models, we introduced random data augmentation during training, including HorizontalFlip, ImageCompression, GaussNoise, GaussianBlur, and Grayscale. The pseudo-code for Pytorch-like is as follows:
1 aug_list = [albumentations.Resize(224, 224)]
2 if random.random() < 0.5:
3 aug_list.append(albumentations.HorizontalFlip(p=1.0))
4 if random.random() < 0.5:
5 quality_score = random.randint(50, 100)
6 aug_list.append(albumentations.ImageCompression(
7 quality_lower=quality_score, quality_upper=quality_score))
8 if random.random() < 0.3:
9 aug_list.append(albumentations.GaussNoise(p=1.0))
10 if random.random() < 0.3:
11 aug_list.append(albumentations.GaussianBlur(blur_limit=(3, 5), p=1.0))
12 if random.random() < 0.001:
13 aug_list.append(albumentations.ToGray(p=1.0))
14 aug_list.append(albumentations.Normalize(mean=(0.485, 0.456, 0.406),
15 std=(0.229, 0.224, 0.225), max_pixel_value=255.0, p=1.0))
16 trans = albumentations.Compose(aug_list)
B.2 Details in degraded video classification task
We applied the following transformations to the videos in $D_{\rm v-ood}$ and utilized models trained on many-to-many tasks for testing. Here, we provide the specific implementation details for each task of degraded video classification:
  • (1)
    H.264 compression: H.264, also known as Advanced Video Coding (AVC), is a widely used video compression standard. In this paper, we set the crf to 28 to compress the video.
  • (2)
    JPEG compression: JPEG compression is a widely used image compression standard designed for efficient compression of digital images. The JPEG algorithm is based on the characteristics of the human visual system, taking advantage of the insensitivity of human eyes to the loss of image details, thus achieving lossy compression of data. In this paper, we set the quality to 35 for the degradation experiment.
  • (3)
    FLIP: We randomly select either Horizontal Flip or Vertical Flip with equal probability for the degradation experiment.
  • (4)
    Crop: We randomly crop the video from the original video with a scale of 71% to 93%.
  • (5)
    Text watermark: We randomly add textual watermarks on the random position of videos.
  • (6)
    Image watermark: We randomly added visual watermarks on the random position of videos.
  • (7)
    Gaussian noise: We add Gaussian blur to the video with a setting of $\sigma=7$.
  • (8)
    Color transform: We randomly select one color transformation from brightness, contrast, saturation, hue and set the parameter to 0.5.
Appendix C Visualization of dataset
As shown in Figure 3-22, we provide visualizations of samples from the dataset. From these figures, it can be seen that our dataset possesses diverse content.
Figure 3: ZeroScope [Wang et al., 2023d] generated samples visualization.
Figure 4: I2VGen-XL [Wang et al., 2023c] generated samples visualization.
Figure 5: SVD [Blattmann et al., 2023b] generated samples visualization example.
Figure 6: VideoCrafter [Chen et al., 2024] generated samples visualization.
Figure 7: Pika [pik, 2022] generated samples visualization.
Figure 8: DynamicCrafter [Xing et al., 2023] generated samples visualization.
Figure 9: SD Zhang et al. [2023b] generated samples visualization.
Figure 10: SEINE [Chen et al., 2023c] generated samples visualization.
Figure 11: Latte [Ma et al., 2024b] generated samples visualization.
Figure 12: OpenSora [Ope, 2024] generated samples visualization.
Figure 13: ModelScope [Wang et al., 2023d] generated samples visualization.
Figure 14: MorphStudio [Mor, 2023] generated samples visualization.
Figure 15: MoonValley [moonvalley.ai, 2022] generated samples visualization.
Figure 16: HotShot [Hot, 2023] generated samples visualization.
Figure 17: Show_1 [Zhang et al., 2023c] generated samples visualization.
Figure 18: Gen2 [Esser et al., 2023] generated samples visualization.
Figure 19: Lavie [Wang et al., 2023a] generated samples visualization.
Figure 20: Sora [Brooks et al., 2024] generated samples visualization.
Figure 21: WildScrape Wei et al. [2023], Feng et al. [2023], Xu et al. [2023c] sample visualization.
Figure 22: Crafter [Chen et al., 2023b] generated samples visualization.

### v3

@@
 A diverse and large-scale collection ensures that the dataset encompasses a wide range of video categories, facilitating detailed evaluations of AI-generated video detection algorithms and their effectiveness across various contexts.
+Real video collection: Considering that fake videos from generators are limited to specific domains determined by training datasets such as Kinetics-400 (Kay et al., 2017b) and Youku-mPLUG (Xu et al., 2023d), we sample parts of videos from those datasets as the real part of the GenVideo dataset.
+Specifically, we randomly sample $953,279$ videos from Youku-mPLUG (Xu et al., 2023d) and randomly slice 10-second segments from each video to form real samples.
 Fake video collection: The guiding principle for collecting fake videos is to ensure maximal diversity in content and generators.
@@
 This pipeline facilitates the production of videos using popular generative mechanisms, including diffusion-based models (Xing et al., 2023; Zhang et al., 2023b; Guo et al., 2023a; Wang et al., 2023c; Blattmann et al., 2023b; Chen et al., 2024; Xing et al., 2023; Zhang et al., 2023b; Chen et al., 2023c), transformer-based methods (Ma et al., 2024b; Ope, 2024) and service-based method Pika website (pik, 2022).
-We selected common categories such as humans, animals, and plants as foreground keywords, and typical scenes like “in the park” or “on the lawn” as background keywords.
-Leveraging a large language model (Le Scao et al., 2023), these foreground and background keywords are expanded into comprehensive textual prompts to guide text-to-video generation.
-For image-to-video generation, we employed various text-to-image models, including different versions of Stable Diffusion (SD (Rombach et al., 2022), SDXL (Podell et al., 2023)).
-Using the enriched textual prompts generated by the large language model, we created corresponding images, which were subsequently used as input for image-to-video models to generate the final videos.
+For image-to-video generation, we employed various text-to-image models to produce diverse images, including different versions of Stable Diffusion (SD (Rombach et al., 2022), SDXL (Podell et al., 2023)).
+In order to produce videos with rich semantics through different generative approaches, including semantic-diverse text and image prompts across person, objects and diverse scenes, we firstly construct a rich prompt dictionary.
+In detail, we selected common 100 categories such as humans, animals, plants and etc as foreground keywords, and typical 20 scenes like “in the park” or “on the lawn” as background keywords.
+Leveraging a large language model (Le Scao et al., 2023), these foreground and background keywords are expanded into about 4000 comprehensive textual prompts.
+Besides, we also randomly sample 1600 textual prompts from VBench Huang et al.
+(2024) with consideration of semantic diversity and style diversity.
+Secondly, each prompt from constructed prompt dictionary is fed into T2V (resp., T2I) generators to produce diverse videos (resp., images).
+I2V generators leverage images generated from T2I generators to produce videos.
 To assemble a representative testing set, we investigated current video generators based on different model architectures (Blattmann et al., 2023a; Brooks et al., 2024; Ghosh et al., 2024) and scraped example videos from their projects, as illustrated by WildScape in Table 1.
@@
 (2023a) that include typical generation methods and demonstrate relatively high generation quality.
-Real video collection: Considering that fake videos from generators are limited to specific domains determined by training datasets such as Kinetics-400 (Kay et al., 2017b) and Youku-mPLUG (Xu et al., 2023d), we sample parts of videos from those datasets as the real part of the GenVideo dataset.
-Specifically, we randomly sample $953,279$ videos from Youku-mPLUG (Xu et al., 2023d) and randomly slice 10-second segments from each video to form real samples.
 4 DeMamba
+Figure 2: The overall framework of our Detail Mamba.
 4.1 Preliminaries
@@
 Finally, we use binary cross-entropy loss to train our model to classify real/fake videos.
-Figure 2: The overall framework of our Detail Mamba (DeMamba).
 5 Experiments
@@
 Table 2: Training parameter settings.
-Task Model Batchsize LR Samples Epochs Optimizer Scheduler Loss
-per epoch
-F3Net 1024 1e-5 $1024\times 2000$ 30 AdamW [20, 25], lr * 0.1 BCE
-NPR 1024 1e-5 $1024\times 2000$ 30 [20, 25], lr * 0.1
-STIL 128 1e-5 $128\times 4000$ 30 [20, 25], lr * 0.1
-VideoMAE-B 128 1e-5 $128\times 4000$ 30 [20, 25], lr * 0.1
-many- CLIP-B-LP 1024 1e-6 $1024\times 2000$ 30 [20, 25], lr * 0.1
-to- CLIP-B-FT 1024 1e-6 $1024\times 2000$ 10 -
-many XCLIP-B-LP 128 1e-6 $128\times 4000$ 30 [20, 25], lr * 0.1
-generalization XCLIP-B-FT 128 1e-6 $128\times 4000$ 10 -
-DeMamba-CLIP-LP 128 1e-6 $1024\times 2000$ 30 [20, 25], lr * 0.1
-DeMamba-CLIP-FT 128 1e-6 $1024\times 2000$ 10 -
-DeMamba-XCLIP-LP 128 1e-6 $128\times 4000$ 30 [20, 25], lr * 0.1
-DeMamba-XCLIP-FT 128 1e-6 $128\times 4000$ 10 -
-one- STIL 128 1e-5 $128\times 400$ 10 AdamW - BCE
-to- NPR 1024 1e-5 $1024\times 2000$ 10 -
-many XCLIP-B-FT 1024 1e-6 $128\times 400$ 10 -
-generalization DeMamba-XCLIP-FT 128 1e-6 $128\times 400$ 10 -
-Table 3: Comparisons to the SOTAs in mean accuracy (ACC) and average precision (AP) on the many-to-many generalization task.
-Model Detection Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real Avg.
+Model Batchsize (Video level) LR Epochs Optimizer Scheduler Loss
+F3Net 128 1e-5 30 AdamW [20, 25], lr * 0.1 BCE
+NPR 128 1e-5 30 [20, 25], lr * 0.1
+STIL 128 1e-5 30 [20, 25], lr * 0.1
+VideoMAE-B 128 1e-5 30 [20, 25], lr * 0.1
+CLIP-B-LP 128 1e-6 30 [20, 25], lr * 0.1
+DeMamba-CLIP-LP 128 1e-6 30 [20, 25], lr * 0.1
+XCLIP-B-LP 128 1e-6 30 [20, 25], lr * 0.1
+DeMamba-XCLIP-LP 128 1e-6 30 [20, 25], lr * 0.1
+CLIP-B-FT 128 1e-6 10 -
+MINTIME-CLIP-B 128 1e-6 10 -
+FTCN-CLIP-B 128 1e-6 10 -
+TALL 128 1e-6 10 -
+XCLIP-B-FT 128 1e-6 10 -
+DeMamba-CLIP-FT 128 1e-6 10 -
+DeMamba-XCLIP-FT 128 1e-6 10 -
+Table 3: Comparisons to the SOTAs in F1 score (F1) and average precision (AP) on the many-to-many generalization task.
+Model Detection Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Avg.
 level Studio Valley Scope Scrape
-F3Net Image ACC 83.93 99.71 98.62 77.57 57.00 36.57 99.52 99.71 89.43 76.78 99.14 83.45
-AP 68.27 99.89 99.67 89.35 85.24 63.17 99.58 99.89 93.80 88.41 - 88.73
-NPR Image ACC 91.07 99.57 99.49 24.29 89.64 57.71 97.12 99.86 94.29 87.80 97.46 85.30
-AP 67.17 99.14 99.20 22.76 93.91 61.76 96.33 99.72 94.15 90.40 - 82.45
-STIL Video ACC 78.57 98.14 98.04 76.00 61.79 53.29 99.36 97.36 94.57 65.01 98.72 83.71
-AP 57.21 99.08 99.32 86.19 82.24 70.43 99.25 98.96 97.18 81.32 - 87.12
-VideoMAE-B Video ACC 67.86 96.00 98.41 96.14 77.14 80.43 97.44 96.93 96.29 68.36 99.71 88.61
-AP 66.49 98.85 99.77 99.27 96.55 95.31 99.49 99.69 99.27 90.74 - 94.54
-CLIP-B-PT Image ACC 85.71 82.43 90.36 71.00 79.29 75.43 89.62 86.29 82.14 75.16 57.22 79.67
-AP 6.78 43.56 70.88 29.97 52.97 35.36 55.52 66.03 44.23 42.99 - 44.83
-DeMamba Video ACC 58.93 96.43 93.12 68.00 69.36 69.00 89.14 91.86 96.14 56.59 98.06 $80.60_{{\color[rgb]{0.453125,0,0}+0.93}}$
--CLIP-PT AP 25.87 95.14 96.23 73.43 83.31 75.49 90.17 95.06 95.05 69.95 - $79.97_{{\color[rgb]{0.453125,0,0}+35.14}}$
-CLIP-B-FT Image ACC 94.64 99.86 91.38 77.29 88.14 86.00 99.68 99.79 84.29 84.67 97.38 91.19
-AP 80.67 99.67 95.24 82.20 93.48 88.62 99.55 99.79 86.93 89.08 - 91.52
-DeMamba Video ACC 95.71 100.00 98.70 69.14 92.43 93.29 100.00 100.00 83.57 82.94 99.44 $92.29_{{\color[rgb]{0.453125,0,0}+1.10}}$
-CLIP-B-FT AP 85.50 100.00 99.59 76.15 96.78 96.99 99.97 100.00 89.80 89.72 - $93.45_{{\color[rgb]{0.453125,0,0}+1.93}}$
-XCLIP-B-PT Video ACC 81.34 82.15 83.35 80.98 81.82 81.55 82.14 82.98 81.93 81.10 81.37 81.88
-AP 16.39 72.16 87.77 39.86 65.57 54.26 75.23 84.80 61.60 55.28 - 61.29
-DeMamba Video ACC 66.07 95.86 94.64 77.86 75.36 80.29 90.89 92.50 96.00 66.41 95.12 $84.64_{{\color[rgb]{0.453125,0,0}+2.76}}$
--XCLIP-PT AP 18.26 93.50 94.72 69.94 78.08 71.50 83.95 92.23 93.54 68.10 - $76.38_{{\color[rgb]{0.453125,0,0}+15.09}}$
-XCLIP-B-FT Video ACC 82.14 99.57 93.62 61.29 79.36 69.71 97.92 99.79 77.14 83.59 98.14 85.66
-AP 64.42 99.73 96.78 70.98 90.35 77.28 97.34 99.84 82.01 88.97 - 86.77
-DeMamba Video ACC 98.21 100.00 99.86 65.43 94.86 98.86 100.00 100.00 92.86 89.09 99.42 $\textbf{94.42}_{{\color[rgb]{0.453125,0,0}+8.76}}$
--XCLIP-FT AP 93.32 100.00 99.97 85.55 98.97 99.60 99.98 100.00 97.77 95.75 - $\textbf{97.10}_{{\color[rgb]{0.453125,0,0}+10.43}}$
+F3Net Image R 0.8393 0.9971 0.9862 0.7757 0.5700 0.3657 0.9952 0.9971 0.8943 0.7678 0.8188
+F1 0.5000 0.9406 0.9628 0.8169 0.6988 0.4904 0.9332 0.9688 0.8873 0.8251 0.8024
+AP 0.6827 0.9989 0.9967 0.8935 0.8524 0.6317 0.9958 0.9989 0.9380 0.8841 0.8873
+NPR Image R 0.9107 0.9957 0.9949 0.2429 0.8964 0.5771 0.9712 0.9986 0.9429 0.8780 0.8408
+F1 0.2786 0.8441 0.9131 0.3028 0.8627 0.5944 0.8170 0.9164 0.8184 0.8163 0.7164
+AP 0.6717 0.9914 0.9920 0.2276 0.9391 0.6176 0.9633 0.9972 0.9415 0.9040 0.8245
+STIL Video R 0.7857 0.9814 0.9804 0.7600 0.6179 0.5329 0.9936 0.9736 0.9457 0.6501 0.8222
+F1 0.3805 0.9068 0.9458 0.7824 0.7232 0.6217 0.9039 0.9433 0.8884 0.7267 0.7823
+AP 0.5721 0.9908 0.9932 0.8619 0.8224 0.7043 0.9925 0.9896 0.9718 0.8132 0.8712
+VideoMAE-B Video R 0.6786 0.9600 0.9841 0.9614 0.7714 0.8043 0.9744 0.9693 0.9629 0.6836 0.8750
+F1 0.623 0.9593 0.9819 0.9600 0.8608 0.8722 0.9644 0.9745 0.9615 0.7977 0.8955
+AP 0.6649 0.9885 0.9977 0.9927 0.9655 0.9531 0.9949 0.9969 0.9927 0.9074 0.9454
+MINTIME-CLIP-B Video R 0.8929 1.0000 0.9899 0.2643 0.9679 0.9814 0.9984 1.0000 0.8429 0.8238 0.8762
+F1 0.4902 0.9340 0.9606 0.3760 0.9499 0.9246 0.9266 0.9659 0.8495 0.8539 0.8231
+AP 0.8321 0.9999 0.9967 0.5084 0.9920 0.9927 0.9976 0.9999 0.9183 0.9177 0.9155
+FTCN-CLIP-B Video R 0.8750 1.0000 0.9891 0.1771 0.9771 0.9186 1.0000 1.0000 0.8529 0.8283 0.8618
+F1 0.7840 0.9859 0.9873 0.2922 0.9813 0.9442 0.9843 0.9929 0.9073 0.8960 0.8755
+AP 0.9179 0.9999 0.9979 0.4594 0.9976 0.9780 0.9999 0.9999 0.9469 0.9232 0.9221
+TALL Video R 0.9107 0.9828 0.9783 0.8300 0.7657 0.7957 0.9952 0.9893 0.9414 0.6631 0.8852
+F1 0.2615 0.8240 0.8964 0.7430 0.7782 0.7234 0.8133 0.9032 0.8027 0.6736 0.7419
+AP 0.7115 0.9689 0.9851 0.7938 0.8459 0.7938 0.9879 0.9902 0.9270 0.7647 0.8791
+CLIP-B-PT Image R 0.8571 0.8243 0.9036 0.7100 0.7929 0.7543 0.8962 0.8629 0.8214 0.7516 0.8174
+F1 0.0215 0.2078 0.1816 0.3271 0.1918 0.2053 0.3509 0.2068 0.2357 0.2289 0.2157
+AP 0.6780 0.4356 0.7088 0.2997 0.5297 0.3536 0.5552 0.6603 0.4423 0.4299 0.4483
+DeMamba-CLIP-PT Video R 0.5893 0.9643 0.9312 0.6800 0.6936 0.6900 0.8914 0.9186 0.9614 0.5659 $0.7886_{{\color[rgb]{0,0.453125,0}-0.0308}}$
+F1 0.234 0.8610 0.8992 0.6949 0.7574 0.7015 0.8105 0.8934 0.8588 0.6379 $0.7349_{{\color[rgb]{0.453125,0,0}+0.5060}}$
+AP 0.2587 0.9514 0.9623 0.7343 0.8331 0.7549 0.9017 0.9506 0.9505 0.6995 $0.7997_{{\color[rgb]{0.453125,0,0}+0.3514}}$
+CLIP-B-FT Image R 0.9464 0.9986 0.9138 0.7729 0.8814 0.8600 0.9968 0.9979 0.8429 0.8467 0.9057
+F1 0.2818 0.8422 0.8691 0.7204 0.8521 0.7698 0.8252 0.9137 0.7608 0.7955 0.7631
+AP 0.8067 0.9967 0.9524 0.8220 0.9348 0.8862 0.9955 0.9979 0.8693 0.8908 0.9152
+DeMamba-CLIP-B-FT Video R 0.9571 1.0000 0.9870 0.6914 0.9243 0.9329 1.0000 1.0000 0.8357 0.8294 $0.9158_{{\color[rgb]{0.453125,0,0}+0.0101}}$
+F1 0.6463 0.9615 0.9739 0.7803 0.9414 0.9276 0.9572 0.9804 0.8723 0.8782 $0.8919_{{\color[rgb]{0.453125,0,0}+0.1288}}$
+AP 0.8550 1.0000 0.9959 0.7615 0.9678 0.9699 0.9997 1.0000 0.8980 0.8972 $0.9345_{{\color[rgb]{0.453125,0,0}+0.1930}}$
+XCLIP-B-PT Video R 0.8134 0.8215 0.8335 0.8098 0.8182 0.8155 0.8214 0.8298 0.8193 0.8110 0.8193
+F1 0.0458 0.3666 0.5237 0.3619 0.5196 0.3640 0.3424 0.5250 0.3656 0.4240 0.3839
+AP 0.1639 0.7216 0.8777 0.3986 0.6557 0.5426 0.7523 0.8480 0.6160 0.5528 0.6129
+DeMamba-XCLIP-PT Video R 0.6607 0.9586 0.9464 0.7786 0.7536 0.8029 0.9089 0.9250 0.9600 0.6641 $0.8359_{{\color[rgb]{0.453125,0,0}+0.0166}}$
+F1 0.1244 0.7223 0.8232 0.6293 0.7172 0.6427 0.6758 0.8137 0.7226 0.6058 $0.6477_{{\color[rgb]{0.453125,0,0}+0.2638}}$
+AP 0.1826 0.9350 0.9472 0.6994 0.7808 0.7150 0.8395 0.9223 0.9354 0.6810 $0.7638_{{\color[rgb]{0.453125,0,0}+0.1509}}$
+XCLIP-B-FT Video R 0.8214 0.9957 0.9362 0.6129 0.7936 0.6971 0.9792 0.9979 0.7714 0.8359 0.8441
+F1 0.3147 0.8805 0.9041 0.6530 0.8242 0.7099 0.8602 0.9370 0.7570 0.8212 0.7662
+AP 0.6442 0.9973 0.9678 0.7098 0.9035 0.7728 0.9734 0.9984 0.8201 0.8897 0.8677
+DeMamba-XCLIP-FT Video R 0.9821 1.0000 0.9986 0.6543 0.9486 0.9886 1.0000 1.0000 0.9286 0.8909 $\textbf{0.9392}_{{\color[rgb]{0.453125,0,0}+0.0951}}$
+F1 0.6467 0.9602 0.9790 0.7539 0.9537 0.9551 0.9557 0.9797 0.9240 0.9120 $\textbf{0.9020}_{{\color[rgb]{0.453125,0,0}+0.1358}}$
+AP 0.9332 1.0000 0.9997 0.8555 0.9897 0.9960 0.9998 1.0000 0.9777 0.9575 $\textbf{0.9710}_{{\color[rgb]{0.453125,0,0}+0.1043}}$
 Table 4: Comparisons to the SOTAs in mean accuracy (ACC) and average precision (AP) on the one-to-many generalization task.
 Training Model Detection Metric Testing subset Avg.
-Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real
+Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild
 subset level Studio Valley Scope Scrape
-Pika NPR Image ACC 55.36 77.57 71.88 4.86 7.21 4.29 86.26 60.29 71.43 31.53 99.52 51.83
-AP 45.74 91.55 92.71 21.80 44.32 22.74 95.04 90.03 84.91 60.88 - 64.97
-STIL Image ACC 75.00 79.43 94.49 57.86 53.14 64.14 97.12 85.29 69.43 62.42 92.43 75.52
-AP 22.35 71.62 93.19 40.61 53.24 47.73 94.94 85.82 58.99 61.91 - 63.04
-XCLIP-B-FT Video ACC 67.86 91.29 96.23 12.00 22.36 9.14 99.84 83.43 75.57 51.84 99.64 64.47
-AP 71.08 97.53 99.44 44.68 72.69 38.37 99.96 97.32 88.00 74.00 - 78.31
-DeMamba Video ACC 92.86 97.29 98.48 38.29 53.50 41.43 99.84 94.07 77.29 64.15 98.65 $\textbf{77.80}_{{\color[rgb]{0.453125,0,0}+13.33}}$
--XCLIP-FT AP 77.75 98.42 99.16 52.97 76.72 56.24 99.80 97.91 82.83 74.81 - $\textbf{81.66}_{{\color[rgb]{0.453125,0,0}+3.35}}$
-SEINE NPR Image ACC 46.43 78.57 63.70 21.86 7.00 3.29 92.97 89.29 33.86 24.84 99.70 51.05
-AP 36.30 92.63 85.02 52.68 25.69 11.05 97.80 97.78 64.64 47.48 - 61.11
-STIL Video ACC 71.43 80.43 88.48 67.71 54.57 55.71 93.93 89.57 72.00 50.11 92.27 74.20
-AP 23.89 71.01 88.18 52.17 54.49 41.23 84.73 87.38 58.72 46.51 - 60.83
-XCLIP-B-FT Video ACC 85.71 95.43 76.23 65.86 35.93 37.00 99.68 99.00 75.57 49.78 99.80 74.54
-AP 85.89 97.97 94.40 92.81 81.68 77.68 98.48 98.91 92.27 67.91 - 88.80
-DeMamba Video ACC 94.64 98.43 92.17 82.43 52.29 54.00 99.52 99.14 79.29 57.88 98.99 $\textbf{82.61}_{{\color[rgb]{0.453125,0,0}+8.07}}$
--XCLIP-FT AP 83.74 99.01 97.66 90.82 84.11 73.30 99.72 99.73 89.72 76.45 - $\textbf{89.43}_{{\color[rgb]{0.453125,0,0}+0.63}}$
-OpenSora NPR Image ACC 55.36 76.29 55.51 58.57 76.50 22.43 74.92 83.07 29.86 60.37 95.95 62.62
-AP 25.65 75.24 65.02 55.12 82.42 20.75 72.65 86.84 28.13 64.50 - 57.63
-STIL Video ACC 32.14 45.43 56.45 35.14 45.07 34.57 57.83 63.14 19.86 43.95 98.13 48.33
-AP 6.94 55.62 75.99 43.55 68.06 44.01 63.84 80.59 29.39 57.58 - 52.56
-XCLIP-B-FT Video ACC 67.86 75.86 67.46 70.86 73.14 43.57 79.87 86.29 33.43 63.17 98.10 69.06
-AP 48.28 81.39 81.84 77.38 86.08 51.87 83.41 93.18 39.27 72.74 - 71.54
-DeMamba Video ACC 55.36 87.43 81.30 73.14 85.21 73.14 89.62 90.07 44.86 58.10 97.30 $\textbf{75.95}_{{\color[rgb]{0.453125,0,0}+6.89}}$
--XCLIP-FT AP 25.89 86.63 87.27 74.38 91.12 76.01 86.41 93.83 48.74 67.92 - $\textbf{73.82}_{{\color[rgb]{0.453125,0,0}+2.28}}$
+Pika NPR Image R 0.5536 0.7757 0.7188 0.4860 0.0721 0.0429 0.8626 0.6029 0.7143 0.3153 0.5144
+F1 0.4627 0.8410 0.8197 0.0871 0.1293 0.0772 0.8894 0.7368 0.8019 0.4604 0.5306
+AP 0.4574 0.9155 0.9271 0.2180 0.4432 0.2274 0.9504 0.9003 0.8491 0.6088 0.6497
+STIL Image R 0.7500 0.7943 0.9449 0.5786 0.5314 0.6414 0.9712 0.8529 0.6943 0.6242 0.7383
+F1 0.0984 0.5530 0.7580 0.4355 0.5128 0.4708 0.6107 0.7130 0.5008 0.5117 0.5165
+AP 0.2235 0.7162 0.9319 0.4061 0.5324 0.4773 0.9494 0.8582 0.5899 0.6191 0.6304
+MINTIME-CLIP-B Video R 0.1607 0.2686 0.1471 0.0557 0.3757 0.0900 0.2188 0.4071 0.0114 0.4309 0.2166
+F1 0.2338 0.4178 0.2545 0.1039 0.5428 0.1626 0.3535 0.5752 0.0222 0.5969 0.3263
+AP 0.1692 0.7433 0.6066 0.3903 0.8229 0.4108 0.7479 0.8862 0.2385 0.7225 0.5738
+FTCN-CLIP-B Video R 0.9107 0.9300 0.9232 0.1771 0.1750 0.0643 0.9968 0.8479 0.7457 0.5173 0.6287
+F1 0.5896 0.9195 0.9371 0.2770 0.2855 0.1088 0.9483 0.8948 0.8106 0.6513 0.6423
+AP 0.8369 0.9705 0.9718 0.3747 0.4490 0.1733 0.9975 0.9541 0.7971 0.6686 0.7194
+TALL Video R 0.6071 0.7043 0.9449 0.5914 0.5757 0.6400 0.9633 0.8871 0.6057 0.6210 0.7141
+F1 0.1038 0.5609 0.8029 0.4932 0.5817 0.5231 0.6726 0.7746 0.5024 0.5569 0.5572
+AP 0.1582 0.6363 0.9347 0.4400 0.5907 0.5106 0.9209 0.8785 0.5111 0.6443 0.6225
+XCLIP-B-FT Video R 0.6786 0.9129 0.9623 0.1200 0.2236 0.0914 0.9984 0.8343 0.7557 0.5184 0.6096
+F1 0.5891 0.9301 0.9679 0.2049 0.3581 0.1579 0.9712 0.8974 0.8361 0.6662 0.6579
+AP 0.7108 0.9753 0.9944 0.4468 0.7269 0.3837 0.9996 0.9732 0.8800 0.7400 0.7831
+DeMamba-XCLIP-FT Video R 0.9286 0.9729 0.9848 0.3829 0.5350 0.4143 0.9984 0.9407 0.7729 0.6415 $\textbf{0.7572}_{{\color[rgb]{0.453125,0,0}+0.1476}}$
+F1 0.4298 0.8990 0.9460 0.4864 0.6559 0.5160 0.9017 0.9235 0.7869 0.7183 $\textbf{0.7263}_{{\color[rgb]{0.453125,0,0}+0.0684}}$
+AP 0.7775 0.9842 0.9916 0.5297 0.7672 0.5624 0.9980 0.9791 0.8283 0.7481 $\textbf{0.8166}_{{\color[rgb]{0.453125,0,0}+0.0335}}$
+SEINE NPR Image R 0.4643 0.7857 0.6370 0.2186 0.0700 0.0329 0.9297 0.8929 0.3386 0.2484 0.4618
+F1 0.4685 0.8592 0.7684 0.3469 0.1283 0.0612 0.9401 0.9332 0.4907 0.3882 0.5385
+AP 0.3630 0.9263 0.8502 0.5268 0.2569 0.1105 0.9780 0.9778 0.6464 0.4748 0.6111
+STIL Video R 0.7143 0.8043 0.8848 0.6771 0.5457 0.5571 0.9393 0.8957 0.7200 0.5011 0.7239
+F1 0.0922 0.5533 0.7240 0.4864 0.5199 0.4181 0.5921 0.7317 0.5099 0.4292 0.5057
+AP 0.2389 0.7101 0.8818 0.5217 0.5449 0.4123 0.8473 0.8738 0.5872 0.4651 0.6083
+MINTIME-CLIP-B Video R 0.6964 0.9257 0.9116 0.1071 0.2293 0.1071 0.9968 0.8600 0.7371 0.5637 0.6135
+F1 0.5532 0.9297 0.9378 0.1827 0.3633 0.1827 0.9630 0.9087 0.8177 0.6988 0.6538
+AP 0.6596 0.9781 0.9790 0.3688 0.6215 0.3311 0.9978 0.9701 0.7963 0.7402 0.7443
+FTCN-CLIP-B Video R 0.8929 0.9786 0.6297 0.4043 0.1650 0.1557 0.9984 0.9971 0.6514 0.4060 0.6279
+F1 0.8130 0.9765 0.7664 0.5660 0.2802 0.2636 0.9858 0.9922 0.7775 0.5697 0.6991
+AP 0.8836 0.9972 0.8736 0.8028 0.6072 0.4321 0.9998 0.9998 0.8139 0.6255 0.8036
+TALL Video R 0.5714 0.7643 0.7957 0.6443 0.3486 0.5629 0.8514 0.9143 0.6714 0.4438 0.6568
+F1 0.1730 0.7053 0.7957 0.6289 0.4498 0.5727 0.7403 0.86423 0.6469 0.5077 0.6085
+AP 0.2716 0.7836 0.8838 0.6896 0.5518 0.6093 0.8240 0.9426 0.7007 0.5483 0.6805
+XCLIP-B-FT Video R 0.8571 0.9543 0.7623 0.6586 0.3593 0.3700 0.9968 0.9900 0.7557 0.4978 0.7201
+F1 0.7705 0.9632 0.8580 0.7814 0.5234 0.5291 0.9826 0.9879 0.8468 0.6548 0.7898
+AP 0.8589 0.9797 0.9440 0.9281 0.8168 0.7768 0.9848 0.9891 0.9227 0.6791 0.8880
+DeMamba-XCLIP-FT Video R 0.9464 0.9843 0.9217 0.8243 0.5229 0.5400 0.9952 0.9914 0.7929 0.5788 $\textbf{0.8098}_{{\color[rgb]{0.453125,0,0}+0.0897}}$
+F1 0.5000 0.9255 0.9240 0.8381 0.6559 0.6412 0.9228 0.9609 0.8192 0.6855 $\textbf{0.7873}_{{\color[rgb]{0,0.453125,0}-0.0025}}$
+AP 0.8374 0.9901 0.9766 0.9082 0.8411 0.7330 0.9972 0.9973 0.8972 0.7645 $\textbf{0.8943}_{{\color[rgb]{0.453125,0,0}+0.0063}}$
+OpenSora NPR Image R 0.5536 0.7629 0.5551 0.5857 0.7650 0.2243 0.7492 0.8307 0.2986 0.6037 0.5929
+F1 0.1263 0.6520 0.6008 0.5406 0.7448 0.2490 0.6248 0.7835 0.3184 0.5918 0.5232
+AP 0.2565 0.7524 0.6502 0.5512 0.8242 0.2075 0.7265 0.8684 0.2813 0.6450 0.5763
+STIL Video R 0.3214 0.4543 0.5645 0.3514 0.4507 0.3457 0.5783 0.6314 0.1986 0.4395 0.4336
+F1 0.1313 0.5282 0.6644 0.4332 0.5686 0.4277 0.6167 0.7153 0.2712 0.5349 0.4892
+AP 0.0694 0.5562 0.7599 0.4355 0.6806 0.4401 0.6384 0.8059 0.2939 0.5758 0.5256
+MINTIME-CLIP-B Video R 0.8571 0.9214 0.7971 0.6343 0.2357 0.3057 0.9712 0.9636 0.7257 0.4773 0.6889
+F1 0.6000 0.9208 0.8675 0.7396 0.3695 0.4412 0.9426 0.9619 0.8044 0.6212 0.7269
+AP 0.7981 0.9739 0.9370 0.8451 0.6232 0.5819 0.9891 0.9926 0.8529 0.6925 0.8286
+FTCN-CLIP-B Video R 0.1964 0.1786 0.1442 0.1957 0.1579 0.0329 0.2716 0.3229 0.0214 0.3413 0.1863
+F1 0.3235 0.3027 0.2519 0.3270 0.2725 0.0635 0.4252 0.4879 0.0419 0.5084 0.3005
+AP 0.3178 0.6675 0.4927 0.6954 0.7089 0.2382 0.7876 0.8500 0.2263 0.7049 0.5689
+TALL Video R 0.3393 0.5771 0.5225 0.3986 0.5893 0.6129 0.5655 0.6736 0.2429 0.3952 0.4916
+F1 0.1357 0.6173 0.6256 0.4713 0.6790 0.6432 0.5975 0.7405 0.3163 0.4890 0.5315
+AP 0.0798 0.6447 0.7246 0.4924 0.7830 0.6916 0.6166 0.8329 0.3132 0.5313 0.5710
+XCLIP-B-FT Video R 0.6786 0.7586 0.6746 0.7086 0.7314 0.4357 0.7987 0.8629 0.3343 0.6317 0.6615
+F1 0.2686 0.7479 0.7443 0.7162 0.7833 0.5096 0.7595 0.8638 0.4167 0.6875 0.6497
+AP 0.4828 0.8139 0.8184 0.7738 0.8608 0.5187 0.8341 0.9318 0.3927 0.7274 0.7154
+DeMamba-XCLIP-FT Video R 0.5536 0.8743 0.8130 0.7314 0.8521 0.7314 0.8962 0.9007 0.4486 0.5810 $\textbf{0.7382}_{{\color[rgb]{0.453125,0,0}+0.0767}}$
+F1 0.1742 0.7742 0.8094 0.6905 0.8333 0.6905 0.7706 0.8604 0.4895 0.6209 $\textbf{0.6713}_{{\color[rgb]{0.453125,0,0}+0.0216}}$
+AP 0.2589 0.8663 0.8727 0.7438 0.9112 0.7601 0.8641 0.9383 0.4874 0.6792 $\textbf{0.7382}_{{\color[rgb]{0.453125,0,0}+0.0228}}$
 Table 5: Robustness evaluation of different detectors on many-to-many generalization task.
@@
 level ${\rm CRF}=28$ JEPG Flip Crop Text Image noise transform
-F3Net Image ACC 83.45 $77.55_{{\color[rgb]{0,0.453125,0}-5.90}}$ $82.04_{{\color[rgb]{0,0.453125,0}-1.41}}$ $81.45_{{\color[rgb]{0,0.453125,0}-2.00}}$ $69.71_{{\color[rgb]{0,0.453125,0}-13.74}}$ $76.90_{{\color[rgb]{0,0.453125,0}-7.55}}$ $78.67_{{\color[rgb]{0,0.453125,0}-4.78}}$ $82.94_{{\color[rgb]{0,0.453125,0}-0.51}}$ $82.93_{{\color[rgb]{0,0.453125,0}-0.52}}$
-AP 88.73 $87.39_{{\color[rgb]{0,0.453125,0}-1.34}}$ $\textbf{81.02}_{{\color[rgb]{0,0.453125,0}-7.71}}$ $87.70_{{\color[rgb]{0,0.453125,0}-1.73}}$ $67.05_{{\color[rgb]{0,0.453125,0}-21.68}}$ $87.01_{{\color[rgb]{0,0.453125,0}-1.72}}$ $87.93_{{\color[rgb]{0,0.453125,0}-0.70}}$ $87.17_{{\color[rgb]{0,0.453125,0}-1.56}}$ $87.17_{{\color[rgb]{0,0.453125,0}-1.56}}$
-NPR Image ACC 85.30 $80.83_{{\color[rgb]{0,0.453125,0}-4.47}}$ $73.92_{{\color[rgb]{0,0.453125,0}-11.38}}$ $83.82_{{\color[rgb]{0,0.453125,0}-1.48}}$ $71.69_{{\color[rgb]{0,0.453125,0}-13.61}}$ $83.26_{{\color[rgb]{0,0.453125,0}-2.04}}$ $82.43_{{\color[rgb]{0,0.453125,0}-2.87}}$ $84.94_{{\color[rgb]{0,0.453125,0}-0.36}}$ $85.03_{{\color[rgb]{0,0.453125,0}-0.47}}$
-AP 82.45 $76.12_{{\color[rgb]{0,0.453125,0}-6.33}}$ $53.72_{{\color[rgb]{0,0.453125,0}-28.73}}$ $79.05_{{\color[rgb]{0,0.453125,0}-3.40}}$ $36.36_{{\color[rgb]{0,0.453125,0}-46.09}}$ $67.16_{{\color[rgb]{0,0.453125,0}-15.28}}$ $77.02_{{\color[rgb]{0,0.453125,0}-5.43}}$ $57.55_{{\color[rgb]{0,0.453125,0}-24.90}}$ $76.81_{{\color[rgb]{0,0.453125,0}-5.64}}$
-STIL Video ACC 83.71 $78.14_{{\color[rgb]{0,0.453125,0}-5.57}}$ $73.79_{{\color[rgb]{0,0.453125,0}-9.92}}$ $81.97_{{\color[rgb]{0,0.453125,0}-1.74}}$ $77.58_{{\color[rgb]{0,0.453125,0}-6.13}}$ $77.64_{{\color[rgb]{0,0.453125,0}-6.07}}$ $74.16_{{\color[rgb]{0,0.453125,0}-9.55}}$ $83.44_{{\color[rgb]{0,0.453125,0}-0.27}}$ $83.47_{{\color[rgb]{0,0.453125,0}-0.24}}$
-AP 87.12 $85.04_{{\color[rgb]{0,0.453125,0}-2.08}}$ $59.38_{{\color[rgb]{0,0.453125,0}-27.74}}$ $85.21_{{\color[rgb]{0,0.453125,0}-1.91}}$ $76.38_{{\color[rgb]{0,0.453125,0}-10.74}}$ $82.59_{{\color[rgb]{0,0.453125,0}-4.53}}$ $83.09_{{\color[rgb]{0,0.453125,0}-4.03}}$ $74.54_{{\color[rgb]{0,0.453125,0}-12.58}}$ $86.65_{{\color[rgb]{0,0.453125,0}-0.47}}$
-CLIP-B-FT Image ACC 91.19 $82.92_{{\color[rgb]{0,0.453125,0}-8.27}}$ $76.78_{{\color[rgb]{0,0.453125,0}-12.41}}$ $89.23_{{\color[rgb]{0,0.453125,0}-12.41}}$ $66.91_{{\color[rgb]{0,0.453125,0}-24.28}}$ $79.65_{{\color[rgb]{0,0.453125,0}-11.54}}$ $78.58_{{\color[rgb]{0,0.453125,0}-12.61}}$ $88.50_{{\color[rgb]{0,0.453125,0}-2.69}}$ $88.81_{{\color[rgb]{0,0.453125,0}-2.38}}$
-AP 91.52 $90.42_{{\color[rgb]{0,0.453125,0}-1.10}}$ $85.33_{{\color[rgb]{0,0.453125,0}-12.41}}$ $91.05_{{\color[rgb]{0,0.453125,0}-0.14}}$ $76.35_{{\color[rgb]{0,0.453125,0}-15.17}}$ $88.94_{{\color[rgb]{0,0.453125,0}-2.58}}$ $91.05_{{\color[rgb]{0,0.453125,0}-0.47}}$ $84.55_{{\color[rgb]{0,0.453125,0}-6.97}}$ $91.05_{{\color[rgb]{0,0.453125,0}-0.50}}$
-XCLIP-B-FT Video ACC 85.66 $84.17_{{\color[rgb]{0,0.453125,0}-1.49}}$ $53.96_{{\color[rgb]{0,0.453125,0}-31.70}}$ $84.48_{{\color[rgb]{0,0.453125,0}-1.18}}$ $56.24_{{\color[rgb]{0,0.453125,0}-30.42}}$ $85.06_{{\color[rgb]{0,0.453125,0}-1.60}}$ $82.84_{{\color[rgb]{0,0.453125,0}-2.82}}$ $82.94_{{\color[rgb]{0,0.453125,0}-2.72}}$ $85.41_{{\color[rgb]{0,0.453125,0}-0.25}}$
-AP 86.77 $86.53_{{\color[rgb]{0,0.453125,0}-0.24}}$ $54.11_{{\color[rgb]{0,0.453125,0}-32.66}}$ $86.76_{{\color[rgb]{0,0.453125,0}-0.01}}$ $46.10_{{\color[rgb]{0,0.453125,0}-40.67}}$ $86.71_{{\color[rgb]{0,0.453125,0}-0.06}}$ $86.74_{{\color[rgb]{0,0.453125,0}-0.03}}$ $68.05_{{\color[rgb]{0,0.453125,0}-14.89}}$ $86.01_{{\color[rgb]{0,0.453125,0}-0.76}}$
-DeMamba Video ACC 92.29 $85.63_{{\color[rgb]{0,0.453125,0}-6.66}}$ $83.49_{{\color[rgb]{0,0.453125,0}-8.80}}$ $90.59_{{\color[rgb]{0,0.453125,0}-1.70}}$ $62.25_{{\color[rgb]{0,0.453125,0}-30.04}}$ $83.34_{{\color[rgb]{0,0.453125,0}-8.95}}$ $87.14_{{\color[rgb]{0,0.453125,0}-5.15}}$ $90.63_{{\color[rgb]{0,0.453125,0}-1.66}}$ $91.50_{{\color[rgb]{0,0.453125,0}-0.79}}$
--CLIP-FT AP 93.45 $90.93_{{\color[rgb]{0,0.453125,0}-2.52}}$ $74.03_{{\color[rgb]{0,0.453125,0}-19.42}}$ $93.44_{{\color[rgb]{0,0.453125,0}-0.01}}$ $61.12_{{\color[rgb]{0,0.453125,0}-32.33}}$ $90.08_{{\color[rgb]{0,0.453125,0}-3.37}}$ $93.35_{{\color[rgb]{0,0.453125,0}-0.10}}$ $82.99_{{\color[rgb]{0,0.453125,0}-10.46}}$ $93.39_{{\color[rgb]{0,0.453125,0}-0.06}}$
-DeMamba Video ACC 94.42 $\textbf{90.52}_{{\color[rgb]{0,0.453125,0}-3.90}}$ $\textbf{93.26}_{{\color[rgb]{0,0.453125,0}-1.16}}$ $\textbf{94.26}_{{\color[rgb]{0,0.453125,0}-0.16}}$ $\textbf{72.98}_{{\color[rgb]{0,0.453125,0}-22.44}}$ $\textbf{89.59}_{{\color[rgb]{0,0.453125,0}-4.83}}$ $\textbf{91.06}_{{\color[rgb]{0,0.453125,0}-3.36}}$ $\textbf{94.36}_{{\color[rgb]{0,0.453125,0}-0.06}}$ $\textbf{93.82}_{{\color[rgb]{0,0.453125,0}-0.60}}$
--XCLIP-FT AP 97.10 $\textbf{96.00}_{{\color[rgb]{0,0.453125,0}-1.10}}$ $76.29_{{\color[rgb]{0,0.453125,0}-20.81}}$ $\textbf{96.76}_{{\color[rgb]{0,0.453125,0}-0.34}}$ $\textbf{67.18}_{{\color[rgb]{0,0.453125,0}-29.92}}$ $\textbf{94.30}_{{\color[rgb]{0,0.453125,0}-2.80}}$ $\textbf{97.07}_{{\color[rgb]{0,0.453125,0}-0.03}}$ $\textbf{85.05}_{{\color[rgb]{0,0.453125,0}-12.05}}$ $\textbf{97.05}_{{\color[rgb]{0,0.453125,0}-0.05}}$
+F3Net Image R 0.8188 $0.7721_{{\color[rgb]{0,0.453125,0}-0.0467}}$ $0.8108_{{\color[rgb]{0,0.453125,0}-0.0080}}$ $0.8041_{{\color[rgb]{0,0.453125,0}-0.0147}}$ $0.6782_{{\color[rgb]{0,0.453125,0}-0.1406}}$ $0.7513_{{\color[rgb]{0,0.453125,0}-0.0675}}$ $0.7668_{{\color[rgb]{0,0.453125,0}-0.0520}}$ $0.8025_{{\color[rgb]{0,0.453125,0}-0.0163}}$ $0.8102_{{\color[rgb]{0,0.453125,0}-0.0086}}$
+F1 0.8024 $0.7967_{{\color[rgb]{0,0.453125,0}-0.0057}}$ $0.6984_{{\color[rgb]{0,0.453125,0}-0.1040}}$ $0.7945_{{\color[rgb]{0,0.453125,0}-0.0079}}$ $0.5988_{{\color[rgb]{0,0.453125,0}-0.2036}}$ $0.7858_{{\color[rgb]{0,0.453125,0}-0.0166}}$ $0.7944_{{\color[rgb]{0,0.453125,0}-0.0080}}$ $0.7910_{{\color[rgb]{0,0.453125,0}-0.0114}}$ $0.7912_{{\color[rgb]{0,0.453125,0}-0.0112}}$
+AP 0.8873 $0.8739_{{\color[rgb]{0,0.453125,0}-0.0134}}$ $0.8102_{{\color[rgb]{0,0.453125,0}-0.0771}}$ $0.8770_{{\color[rgb]{0,0.453125,0}-0.0173}}$ $0.6705_{{\color[rgb]{0,0.453125,0}-0.2168}}$ $0.8701_{{\color[rgb]{0,0.453125,0}-0.0172}}$ $0.8793_{{\color[rgb]{0,0.453125,0}-0.0070}}$ $0.8717_{{\color[rgb]{0,0.453125,0}-0.0156}}$ $0.8717_{{\color[rgb]{0,0.453125,0}-0.0156}}$
+NPR Image R 0.8408 $0.8012_{{\color[rgb]{0,0.453125,0}-0.0396}}$ $0.7375_{{\color[rgb]{0,0.453125,0}-0.1033}}$ $0.8365_{{\color[rgb]{0,0.453125,0}-0.0043}}$ $0.7144_{{\color[rgb]{0,0.453125,0}-0.1264}}$ $0.8285_{{\color[rgb]{0,0.453125,0}-0.0123}}$ $0.8200_{{\color[rgb]{0,0.453125,0}-0.0208}}$ $0.8392_{{\color[rgb]{0,0.453125,0}-0.0016}}$ $0.8378_{{\color[rgb]{0,0.453125,0}-0.0030}}$
+F1 0.7164 $0.6382_{{\color[rgb]{0,0.453125,0}-0.0782}}$ $0.4588_{{\color[rgb]{0,0.453125,0}-0.2576}}$ $0.6896_{{\color[rgb]{0,0.453125,0}-0.0268}}$ $0.1655_{{\color[rgb]{0,0.453125,0}-0.5509}}$ $0.5580_{{\color[rgb]{0,0.453125,0}-0.1584}}$ $0.6583_{{\color[rgb]{0,0.453125,0}-0.0581}}$ $0.4791_{{\color[rgb]{0,0.453125,0}-0.2373}}$ $0.6564_{{\color[rgb]{0,0.453125,0}-0.0600}}$
+AP 0.8245 $0.7612_{{\color[rgb]{0,0.453125,0}-0.0633}}$ $0.5372_{{\color[rgb]{0,0.453125,0}-0.2873}}$ $0.7905_{{\color[rgb]{0,0.453125,0}-0.0340}}$ $0.3636_{{\color[rgb]{0,0.453125,0}-0.4609}}$ $0.6716_{{\color[rgb]{0,0.453125,0}-0.1528}}$ $0.7702_{{\color[rgb]{0,0.453125,0}-0.0543}}$ $0.5755_{{\color[rgb]{0,0.453125,0}-0.2490}}$ $0.7681_{{\color[rgb]{0,0.453125,0}-0.0564}}$
+STIL Video R 0.8222 $0.7712_{{\color[rgb]{0,0.453125,0}-0.0510}}$ $0.7215_{{\color[rgb]{0,0.453125,0}-0.1007}}$ $0.8076_{{\color[rgb]{0,0.453125,0}-0.0146}}$ $0.7678_{{\color[rgb]{0,0.453125,0}-0.0544}}$ $0.7694_{{\color[rgb]{0,0.453125,0}-0.0528}}$ $0.7401_{{\color[rgb]{0,0.453125,0}-0.0821}}$ $0.8205_{{\color[rgb]{0,0.453125,0}-0.0017}}$ $0.8195_{{\color[rgb]{0,0.453125,0}-0.0027}}$
+F1 0.7823 $0.7412_{{\color[rgb]{0,0.453125,0}-0.0411}}$ $0.5124_{{\color[rgb]{0,0.453125,0}-0.2699}}$ $0.7435_{{\color[rgb]{0,0.453125,0}-0.0388}}$ $0.6589_{{\color[rgb]{0,0.453125,0}-0.1234}}$ $0.7276_{{\color[rgb]{0,0.453125,0}-0.0547}}$ $0.7301_{{\color[rgb]{0,0.453125,0}-0.0522}}$ $0.6472_{{\color[rgb]{0,0.453125,0}-0.1351}}$ $0.7566_{{\color[rgb]{0,0.453125,0}-0.0257}}$
+AP 0.8712 $0.8504_{{\color[rgb]{0,0.453125,0}-0.0208}}$ $0.5938_{{\color[rgb]{0,0.453125,0}-0.2774}}$ $0.8521_{{\color[rgb]{0,0.453125,0}-0.0191}}$ $0.7638_{{\color[rgb]{0,0.453125,0}-0.1074}}$ $0.8259_{{\color[rgb]{0,0.453125,0}-0.0453}}$ $0.8309_{{\color[rgb]{0,0.453125,0}-0.0403}}$ $0.7454_{{\color[rgb]{0,0.453125,0}-0.1258}}$ $0.8665_{{\color[rgb]{0,0.453125,0}-0.0047}}$
+MINTIME-CLIP-B Video R 0.8724 $0.8215_{{\color[rgb]{0,0.453125,0}-0.0509}}$ $0.8199_{{\color[rgb]{0,0.453125,0}-0.0525}}$ $0.8215_{{\color[rgb]{0,0.453125,0}-0.0508}}$ $0.5534_{{\color[rgb]{0,0.453125,0}-0.3190}}$ $0.8084_{{\color[rgb]{0,0.453125,0}-0.0640}}$ $0.7972_{{\color[rgb]{0,0.453125,0}-0.0752}}$ $0.8628_{{\color[rgb]{0,0.453125,0}-0.0096}}$ $0.8715_{{\color[rgb]{0,0.453125,0}-0.0009}}$
+F1 0.8700 $0.8486_{{\color[rgb]{0,0.453125,0}-0.0214}}$ $\textbf{0.8417}_{{\color[rgb]{0,0.453125,0}-0.0283}}$ $0.8486_{{\color[rgb]{0,0.453125,0}-0.0214}}$ $0.5485_{{\color[rgb]{0,0.453125,0}-0.3215}}$ $0.8468_{{\color[rgb]{0,0.453125,0}-0.0232}}$ $0.8379_{{\color[rgb]{0,0.453125,0}-0.0321}}$ $0.4734_{{\color[rgb]{0,0.453125,0}-0.3966}}$ $0.8658_{{\color[rgb]{0,0.453125,0}-0.0042}}$
+AP 0.9369 $0.9207_{{\color[rgb]{0,0.453125,0}-0.0158}}$ $\textbf{0.9143}_{{\color[rgb]{0,0.453125,0}-0.0226}}$ $0.9207_{{\color[rgb]{0,0.453125,0}-0.0158}}$ $0.6199_{{\color[rgb]{0,0.453125,0}-0.3170}}$ $0.9071_{{\color[rgb]{0,0.453125,0}-0.0298}}$ $0.9227_{{\color[rgb]{0,0.453125,0}-0.0142}}$ $0.8146_{{\color[rgb]{0,0.453125,0}-0.1223}}$ $0.9364_{{\color[rgb]{0,0.453125,0}-0.0005}}$
+FTCN-CLIP-B Video R 0.8935 $0.8519_{{\color[rgb]{0,0.453125,0}-0.0416}}$ $0.8127_{{\color[rgb]{0,0.453125,0}-0.0808}}$ $0.8858_{{\color[rgb]{0,0.453125,0}-0.0077}}$ $0.5546_{{\color[rgb]{0,0.453125,0}-0.3389}}$ $0.8284_{{\color[rgb]{0,0.453125,0}-0.0651}}$ $0.8547_{{\color[rgb]{0,0.453125,0}-0.0388}}$ $0.8874_{{\color[rgb]{0,0.453125,0}-0.0061}}$ $0.8904_{{\color[rgb]{0,0.453125,0}-0.0031}}$
+F1 0.8739 $0.8730_{{\color[rgb]{0,0.453125,0}-0.0009}}$ $0.5495_{{\color[rgb]{0,0.453125,0}-0.3244}}$ $0.8484_{{\color[rgb]{0,0.453125,0}-0.0451}}$ $0.5573_{{\color[rgb]{0,0.453125,0}-0.3166}}$ $0.8689_{{\color[rgb]{0,0.453125,0}-0.0050}}$ $0.8648_{{\color[rgb]{0,0.453125,0}-0.0091}}$ $0.6002_{{\color[rgb]{0,0.453125,0}-0.2735}}$ $0.8638_{{\color[rgb]{0,0.453125,0}-0.0101}}$
+AP 0.9418 $0.9294_{{\color[rgb]{0,0.453125,0}-0.0124}}$ $0.6884_{{\color[rgb]{0,0.453125,0}-0.2534}}$ $0.9316_{{\color[rgb]{0,0.453125,0}-0.0102}}$ $0.6015_{{\color[rgb]{0,0.453125,0}-0.3403}}$ $0.9256_{{\color[rgb]{0,0.453125,0}-0.0162}}$ $0.9386_{{\color[rgb]{0,0.453125,0}-0.0032}}$ $0.8503_{{\color[rgb]{0,0.453125,0}-0.0915}}$ $0.9389_{{\color[rgb]{0,0.453125,0}-0.0029}}$
+TALL Video R 0.8852 $0.8413_{{\color[rgb]{0,0.453125,0}-0.0439}}$ $0.4272_{{\color[rgb]{0,0.453125,0}-0.4580}}$ $0.8785_{{\color[rgb]{0,0.453125,0}-0.0067}}$ $0.7233_{{\color[rgb]{0,0.453125,0}-0.1619}}$ $0.8061_{{\color[rgb]{0,0.453125,0}-0.0791}}$ $0.8045_{{\color[rgb]{0,0.453125,0}-0.0807}}$ $0.8834_{{\color[rgb]{0,0.453125,0}-0.0018}}$ $0.8816_{{\color[rgb]{0,0.453125,0}-0.0036}}$
+F1 0.7419 $0.7411_{{\color[rgb]{0,0.453125,0}-0.0008}}$ $0.5309_{{\color[rgb]{0,0.453125,0}-0.2110}}$ $0.7248_{{\color[rgb]{0,0.453125,0}-0.0171}}$ $0.5844_{{\color[rgb]{0,0.453125,0}-0.1575}}$ $0.7329_{{\color[rgb]{0,0.453125,0}-0.0090}}$ $0.7360_{{\color[rgb]{0,0.453125,0}-0.0059}}$ $0.5996_{{\color[rgb]{0,0.453125,0}-0.1423}}$ $0.6719_{{\color[rgb]{0,0.453125,0}-0.0700}}$
+AP 0.8791 $0.8673_{{\color[rgb]{0,0.453125,0}-0.0118}}$ $0.6439_{{\color[rgb]{0,0.453125,0}-0.2352}}$ $0.8774_{{\color[rgb]{0,0.453125,0}-0.0017}}$ $0.6744_{{\color[rgb]{0,0.453125,0}-0.2047}}$ $0.8608_{{\color[rgb]{0,0.453125,0}-0.0183}}$ $0.8767_{{\color[rgb]{0,0.453125,0}-0.0024}}$ $0.7880_{{\color[rgb]{0,0.453125,0}-0.0911}}$ $0.8439_{{\color[rgb]{0,0.453125,0}-0.0352}}$
+CLIP-B-FT Image R 0.9057 $0.8253_{{\color[rgb]{0,0.453125,0}-0.0804}}$ $0.7632_{{\color[rgb]{0,0.453125,0}-0.1425}}$ $0.8878_{{\color[rgb]{0,0.453125,0}-0.0179}}$ $0.6782_{{\color[rgb]{0,0.453125,0}-0.2275}}$ $0.7877_{{\color[rgb]{0,0.453125,0}-0.1180}}$ $0.7832_{{\color[rgb]{0,0.453125,0}-0.1225}}$ $0.8812_{{\color[rgb]{0,0.453125,0}-0.0245}}$ $0.8809_{{\color[rgb]{0,0.453125,0}-0.0248}}$
+F1 0.7631 $0.7476_{{\color[rgb]{0,0.453125,0}-0.0155}}$ $0.6982_{{\color[rgb]{0,0.453125,0}-0.0649}}$ $0.7533_{{\color[rgb]{0,0.453125,0}-0.0098}}$ $0.6273_{{\color[rgb]{0,0.453125,0}-0.1358}}$ $0.7125_{{\color[rgb]{0,0.453125,0}-0.0506}}$ $0.7537_{{\color[rgb]{0,0.453125,0}-0.0094}}$ $0.6743_{{\color[rgb]{0,0.453125,0}-0.0888}}$ $0.7542_{{\color[rgb]{0,0.453125,0}-0.0089}}$
+AP 0.9152 $0.9042_{{\color[rgb]{0,0.453125,0}-0.0110}}$ $0.8533_{{\color[rgb]{0,0.453125,0}-0.0619}}$ $0.9105_{{\color[rgb]{0,0.453125,0}-0.0047}}$ $0.7635_{{\color[rgb]{0,0.453125,0}-0.1517}}$ $0.8894_{{\color[rgb]{0,0.453125,0}-0.0258}}$ $0.9105_{{\color[rgb]{0,0.453125,0}-0.0047}}$ $0.8455_{{\color[rgb]{0,0.453125,0}-0.0697}}$ $0.9105_{{\color[rgb]{0,0.453125,0}-0.0047}}$
+DeMamba-CLIP-FT Video R 0.9158 $0.8572_{{\color[rgb]{0,0.453125,0}-0.0586}}$ $0.8479_{{\color[rgb]{0,0.453125,0}-0.0679}}$ $0.9047_{{\color[rgb]{0,0.453125,0}-0.0111}}$ $0.6172_{{\color[rgb]{0,0.453125,0}-0.2986}}$ $0.8352_{{\color[rgb]{0,0.453125,0}-0.0806}}$ $0.8664_{{\color[rgb]{0,0.453125,0}-0.0494}}$ $0.9037_{{\color[rgb]{0,0.453125,0}-0.0121}}$ $0.9105_{{\color[rgb]{0,0.453125,0}-0.0053}}$
+F1 0.8919 $0.8750_{{\color[rgb]{0,0.453125,0}-0.0169}}$ $0.6325_{{\color[rgb]{0,0.453125,0}-0.2594}}$ $0.8912_{{\color[rgb]{0,0.453125,0}-0.0007}}$ $0.5233_{{\color[rgb]{0,0.453125,0}-0.3686}}$ $0.8455_{{\color[rgb]{0,0.453125,0}-0.0464}}$ $0.8895_{{\color[rgb]{0,0.453125,0}-0.0024}}$ $0.7672_{{\color[rgb]{0,0.453125,0}-0.1247}}$ $0.8910_{{\color[rgb]{0,0.453125,0}-0.0009}}$
+AP 0.9345 $0.9093_{{\color[rgb]{0,0.453125,0}-0.0252}}$ $0.7403_{{\color[rgb]{0,0.453125,0}-0.1942}}$ $0.9344_{{\color[rgb]{0,0.453125,0}-0.0001}}$ $0.6112_{{\color[rgb]{0,0.453125,0}-0.3233}}$ $0.9008_{{\color[rgb]{0,0.453125,0}-337}}$ $0.9335_{{\color[rgb]{0,0.453125,0}-0.0010}}$ $0.8299_{{\color[rgb]{0,0.453125,0}-0.1046}}$ $0.9339_{{\color[rgb]{0,0.453125,0}-0.0006}}$
+XCLIP-B-FT Video R 0.8441 $0.8387_{{\color[rgb]{0,0.453125,0}-0.0054}}$ $0.5377_{{\color[rgb]{0,0.453125,0}-0.3064}}$ $0.8378_{{\color[rgb]{0,0.453125,0}-0.0063}}$ $0.5597_{{\color[rgb]{0,0.453125,0}-0.2844}}$ $0.8388_{{\color[rgb]{0,0.453125,0}-0.0023}}$ $0.8254_{{\color[rgb]{0,0.453125,0}-0.0187}}$ $0.8195_{{\color[rgb]{0,0.453125,0}-0.0216}}$ $0.8421_{{\color[rgb]{0,0.453125,0}-0.0020}}$
+F1 0.7662 $0.7329_{{\color[rgb]{0,0.453125,0}-0.0333}}$ $0.4188_{{\color[rgb]{0,0.453125,0}-0.3474}}$ $0.7592_{{\color[rgb]{0,0.453125,0}-0.0070}}$ $0.3062_{{\color[rgb]{0,0.453125,0}-0.4600}}$ $0.7595_{{\color[rgb]{0,0.453125,0}-0.0067}}$ $0.7593_{{\color[rgb]{0,0.453125,0}-0.0069}}$ $0.6322_{{\color[rgb]{0,0.453125,0}-0.1340}}$ $0.7488_{{\color[rgb]{0,0.453125,0}-0.0174}}$
+AP 0.8677 $0.8653_{{\color[rgb]{0,0.453125,0}-0.0024}}$ $0.5411_{{\color[rgb]{0,0.453125,0}-0.3266}}$ $0.8676_{{\color[rgb]{0,0.453125,0}-0.0001}}$ $0.4610_{{\color[rgb]{0,0.453125,0}-0.4067}}$ $0.8671_{{\color[rgb]{0,0.453125,0}-0.0006}}$ $0.8674_{{\color[rgb]{0,0.453125,0}-0.0003}}$ $0.6805_{{\color[rgb]{0,0.453125,0}-0.1489}}$ $0.8601_{{\color[rgb]{0,0.453125,0}-0.0076}}$
+DeMamba-XCLIP-FT Video R 0.9392 $\textbf{0.9052}_{{\color[rgb]{0,0.453125,0}-0.0340}}$ $\textbf{0.9326}_{{\color[rgb]{0,0.453125,0}-0.0116}}$ $\textbf{0.9376}_{{\color[rgb]{0,0.453125,0}-0.0016}}$ $\textbf{0.7298}_{{\color[rgb]{0,0.453125,0}-0.2244}}$ $\textbf{0.8915}_{{\color[rgb]{0,0.453125,0}-0.0477}}$ $\textbf{0.9094}_{{\color[rgb]{0,0.453125,0}-0.0288}}$ $\textbf{0.9382}_{{\color[rgb]{0,0.453125,0}-0.0010}}$ $\textbf{0.9342}_{{\color[rgb]{0,0.453125,0}-0.0050}}$
+F1 0.9020 $\textbf{0.8770}_{{\color[rgb]{0,0.453125,0}-0.0250}}$ $0.6823_{{\color[rgb]{0,0.453125,0}-0.2197}}$ $\textbf{0.8820}_{{\color[rgb]{0,0.453125,0}-0.0200}}$ $\textbf{0.6595}_{{\color[rgb]{0,0.453125,0}-0.2425}}$ $\textbf{0.8755}_{{\color[rgb]{0,0.453125,0}-0.0265}}$ $\textbf{0.8990}_{{\color[rgb]{0,0.453125,0}-0.0030}}$ $\textbf{0.7962}_{{\color[rgb]{0,0.453125,0}-0.1058}}$ $\textbf{0.9005}_{{\color[rgb]{0,0.453125,0}-0.0005}}$
+AP 0.9710 $\textbf{0.9600}_{{\color[rgb]{0,0.453125,0}-0.0110}}$ $0.7629_{{\color[rgb]{0,0.453125,0}-0.2081}}$ $\textbf{0.9676}_{{\color[rgb]{0,0.453125,0}-0.0034}}$ $\textbf{0.6718}_{{\color[rgb]{0,0.453125,0}-0.2992}}$ $\textbf{0.9430}_{{\color[rgb]{0,0.453125,0}-0.0280}}$ $\textbf{0.9707}_{{\color[rgb]{0,0.453125,0}-0.0003}}$ $\textbf{0.8505}_{{\color[rgb]{0,0.453125,0}-0.1205}}$ $\textbf{0.9705}_{{\color[rgb]{0,0.453125,0}-0.0005}}$
 Table 6: Ablation study of DeMamba.
-Model Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real Avg.
+Model Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Avg.
 Studio Valley Scope Scrape
-w/o Demamba ACC 82.14 99.57 93.62 61.29 79.36 69.71 97.92 99.79 77.14 83.59 98.14 $85.66_{{\color[rgb]{0,0.453125,0}-8.76}}$
-Block AP 64.42 99.73 96.78 70.98 90.35 77.28 97.34 99.84 82.01 88.97 - $86.77_{{\color[rgb]{0,0.453125,0}-10.43}}$
-w/o global ACC 97.33 99.85 99.45 67.66 93.55 97.88 100.00 100.00 91.34 86.54 98.55 $93.83_{{\color[rgb]{0,0.453125,0}-0.59}}$
-feat AP 87.32 99.35 99.22 82.44 97.99 99.42 99.42 99.85 97.32 94.98 - $95.73_{{\color[rgb]{0,0.453125,0}-1.37}}$
-DeMamba ACC 98.21 100.00 99.86 65.43 94.86 98.86 100.00 100.00 92.86 89.09 99.42 94.42
--XCLIP-FT AP 93.32 100.00 99.97 85.55 98.97 99.60 99.98 100.00 97.77 95.75 - 97.10
+w/o Demamba R 0.8214 0.9957 0.9362 0.6129 0.7936 0.6971 0.9792 0.9979 0.7714 0.8359 $0.8411_{{\color[rgb]{0,0.453125,0}-0.0971}}$
+F1 0.3147 0.8805 0.9041 0.6530 0.8242 0.7099 0.8602 0.9370 0.7570 0.8212 $0.7662_{{\color[rgb]{0,0.453125,0}-0.1358}}$
+AP 0.6442 0.9973 0.9678 0.7098 0.9035 0.7728 0.9734 0.9984 0.8201 0.8897 $0.8677_{{\color[rgb]{0,0.453125,0}-0.1043}}$
+w/o global R 0.9733 0.9985 0.9945 0.6766 0.9355 0.9788 1.0000 1.0000 0.9134 0.8654 $0.9336_{{\color[rgb]{0,0.453125,0}-0.0046}}$
+F1 0.4252 0.9053 0.9475 0.7183 0.9176 0.8960 0.8962 0.9508 0.8618 0.8562 $0.8375_{{\color[rgb]{0,0.453125,0}-0.0645}}$
+AP 0.8732 0.9935 0.9922 0.8244 0.9799 0.9942 0.9942 0.9985 0.9732 0.9498 $0.9573_{{\color[rgb]{0,0.453125,0}-0.0137}}$
+DeMamba-XCLIP-FT R 0.9821 1.0000 0.9986 0.6543 0.9486 0.9886 1.0000 1.0000 0.9286 0.8909 0.9382
+F1 0.6467 0.9602 0.9790 0.7539 0.9537 0.9551 0.9557 0.9797 0.9240 0.9120 0.9020
+AP 0.9332 1.0000 0.9997 0.8555 0.9897 0.9960 0.9998 1.0000 0.9777 0.9575 0.9710
 Table 7: Influence of different zone sizes in DeMamba.
-Model Zone Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real Avg.
+Model Zone Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Avg.
 size Studio Valley Scope Scrape
-$1\times 1$ ACC 100.00 99.86 84.78 95.57 81.43 82.29 100.00 99.79 78.29 81.10 97.64 90.98
-AP 93.80 99.64 91.78 96.03 90.04 86.13 99.92 99.81 82.12 86.96 - 92.62
-$2\times 2$ ACC 95.71 100.00 98.70 69.14 92.43 93.29 100.00 100.00 83.57 82.94 99.44 92.29
-DeMamba AP 85.50 100.00 99.59 76.15 96.78 96.99 99.97 100.00 89.80 89.72 - 93.45
--CLIP-FT $7\times 7$ ACC 98.21 100.00 98.41 65.00 94.00 95.43 100.00 100.00 80.86 84.34 98.19 91.62
-AP 87.28 99.99 99.06 73.51 97.47 97.00 99.76 99.97 85.75 89.95 - 92.97
-$14\times 14$ ACC 100.00 99.86 93.41 88.57 81.71 70.14 100.00 99.93 76.43 80.02 98.70 89.89
-AP 94.94 99.42 97.03 93.56 94.02 84.56 99.12 99.22 77.60 89.46 - 92.89
-$1\times 1$ ACC 91.07 100.00 99.35 73.71 89.50 96.29 100.00 100.00 94.71 81.21 99.37 93.20
-AP 86.65 99.99 99.83 90.64 97.85 98.81 99.91 99.99 98.43 92.18 - 96.43
-$2\times 2$ ACC 98.21 100.00 99.86 65.43 94.86 98.86 100.00 100.00 92.86 89.09 99.42 94.42
-DeMamba AP 93.32 100.00 99.97 85.55 98.97 99.60 99.98 100.00 97.77 95.75 - 97.10
--XCLIP-FT $7\times 7$ ACC 96.43 100.00 99.57 56.00 88.71 96.00 100.00 100.00 94.43 79.48 99.64 91.84
-AP 96.02 100.00 99.97 82.85 99.69 99.52 99.98 100.00 86.51 97.45 - 96.20
-$14\times 14$ ACC 76.79 100.00 99.06 22.86 84.71 95.14 99.52 100.00 87.71 72.25 99.87 85.26
-AP 80.43 100.00 99.76 65.69 98.49 99.31 99.79 99.91 96.87 88.31 - 92.86
+$1\times 1$ R 1.0000 0.9986 0.8478 0.9557 0.8143 0.8229 1.0000 0.9979 0.7829 0.8110 0.9031
+F1 0.3218 0.8556 0.8398 0.8334 0.8216 0.7624 0.8414 0.9215 0.7390 0.7849 0.7722
+AP 0.9380 0.9964 0.9178 0.9603 0.9004 0.8613 0.9992 0.9981 0.8212 0.8696 0.9262
+$2\times 2$ R 0.9571 1.0000 0.9870 0.6914 0.9243 0.9329 1.0000 1.0000 0.8357 0.8294 0.9158
+F1 0.6463 0.9615 0.9739 0.7803 0.9414 0.9276 0.9572 0.9804 0.8723 0.8782 0.8919
+DeMamba AP 0.8550 1.0000 0.9959 0.7615 0.9678 0.9699 0.9997 1.0000 0.8980 0.8972 0.9345
+-CLIP-FT $7\times 7$ R 0.9821 1.0000 0.9841 0.6500 0.9400 0.9543 1.0000 1.0000 0.8086 0.8434 0.9163
+F1 0.3724 0.8855 0.9308 0.6811 0.9085 0.8630 0.8737 0.9393 0.7828 0.8271 0.8064
+AP 0.8728 0.9999 0.9906 0.7351 0.9747 0.9700 0.9976 0.9997 0.8575 0.8995 0.9297
+$14\times 14$ R 1.0000 0.9986 0.9341 0.8857 0.8171 0.7014 1.0000 0.9993 0.7643 0.8002 0.8901
+F1 0.4628 0.9149 0.9214 0.8550 0.8555 0.7430 0.9059 0.9556 0.7845 0.8245 0.8223
+AP 0.9494 0.9942 0.9703 0.9356 0.9402 0.8456 0.9912 0.9922 0.7760 0.8946 0.9289
+$1\times 1$ R 0.9107 1.0000 0.9935 0.7371 0.8950 0.9629 1.0000 1.0000 0.9471 0.8121 0.9258
+F1 0.5952 0.9569 0.9748 0.8066 0.9227 0.9387 0.9521 0.9780 0.9298 0.8644 0.8919
+AP 0.8665 0.9999 0.9983 0.9064 0.9785 0.9881 0.9991 0.9999 0.9843 0.9218 0.9643
+$2\times 2$ R 0.9821 1.0000 0.9986 0.6543 0.9486 0.9886 1.0000 1.0000 0.9286 0.8909 0.9392
+F1 0.6467 0.9602 0.9790 0.7539 0.9537 0.9551 0.9557 0.9797 0.9240 0.9120 0.9020
+DeMamba AP 0.9332 1.0000 0.9997 0.8555 0.9897 0.9960 0.9998 1.0000 0.9777 0.9575 0.9710
+-XCLIP-FT $7\times 7$ R 0.9643 1.0000 0.9957 0.5600 0.8871 0.9600 1.0000 1.0000 0.9443 0.7948 0.9106
+F1 0.7013 0.9689 0.9821 0.6895 0.9244 0.9485 0.9653 0.9842 0.9409 0.8622 0.8967
+AP 0.9602 1.0000 0.9997 0.8285 0.9969 0.9952 0.9998 1.0000 0.8651 0.9745 0.9620
+$14\times 14$ R 0.7679 1.0000 0.9906 0.2286 0.8471 0.9514 0.9952 1.0000 0.8771 0.7225 0.8380
+F1 0.7748 0.9908 0.9909 0.3670 0.9126 0.9659 0.9873 0.9954 0.9253 0.8326 0.8742
+AP 0.8043 1.0000 0.9976 0.6569 0.9849 0.9931 0.9979 0.9991 0.9687 0.8831 0.9286
 Table 8: Influence of scanning order in DeMamba.
-Model Scan Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real Avg.
+Model Scan Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Avg.
 Rrder Studio Valley Scope Scrape
-sweep ACC 95.21 100.00 92.55 61.77 92.69 96.33 99.87 99.32 81.53 73.11 98.79 $90.11_{{\color[rgb]{0,0.453125,0}-3.31}}$
-DeMamba AP 93.32 99.95 97.32 85.55 95.88 97.54 99.66 97.54 87.13 80.82 - $93.47_{{\color[rgb]{0,0.453125,0}-3.63}}$
--XCLIP-FT continuous ACC 98.21 100.00 99.86 65.43 94.86 98.86 100.00 100.00 92.86 89.09 99.42 94.42
-AP 93.32 100.00 99.97 85.55 98.97 99.60 99.98 100.00 97.77 95.75 - 97.10
+sweep R 0.9521 1.0000 0.9255 0.6177 0.9269 0.9633 0.9987 0.9932 0.8153 0.7311 0.8924
+F1 0.4649 0.9211 0.9200 0.6906 0.9212 0.9029 0.9124 0.9557 0.8207 0.7856 $0.8295_{{\color[rgb]{0,0.453125,0}-0.0725}}$
+DeMamba AP 0.9332 0.9995 0.9732 0.8555 0.9588 0.9754 0.9966 0.9754 0.8713 0.8082 $0.9347_{{\color[rgb]{0,0.453125,0}-0.0363}}$
+-XCLIP-FT continuous R 0.9821 1.0000 0.9986 0.6543 0.9486 0.9886 1.0000 1.0000 0.9286 0.8909 0.9392
+F1 0.6467 0.9602 0.9790 0.7539 0.9537 0.9551 0.9557 0.9797 0.9240 0.9120 0.9020
+AP 0.9332 1.0000 0.9997 0.8555 0.9897 0.9960 0.9998 1.0000 0.9777 0.9575 0.9710
 Evaluation metrics.
@@
 Baselines.
-1) CLIP (ICML’21) (Radford et al., 2021) is an innovative model that connects images with textual descriptions, enabling it to perform exceptionally well across various vision and language tasks.
-2) F3Net (ECCV’20) (Qian et al., 2020) addresses the detection of sophisticated face manipulations using a combination of frequency-aware components and loc al statistics.
-Its dual-stream approach effectively identifies forgery patterns.
-3) VideoMAE (NeurIPS’22) (Tong et al., 2022) adapts the Masked Autoencoder (He et al., 2022) concept for video, enhancing feature extraction and improving performance on downstream tasks through self-supervised training.
-4) XCLIP (ECCV’22) (Ni et al., 2022) adapts existing image-language models for video recognition, adding a new cross-frame attention mechanism to improve the exchange of temporal information and tailor prompts specifically for videos.
-5) STIL (MM’21) (Gu et al., 2021a) proposes a novel spatial-temporal inconsistency learning method, which captures the spatial-temporal inconsistencies in forged videos by introducing a spatial inconsistency module, a temporal inconsistency module, and an information supplementation module.
-6) NPR (CVPR’24) (Tan et al., 2024) is a concept introduced to understand and describe the complex, local interdependencies between image pixels that arise from upsampling processes, particularly evident in synthetic images created by generative models like GANs and diffusion models, which are used to identify and analyze the structural artifacts resulting from these operations.
+Three type of image-level methods and six video-level methods are treated as baselines for comparison: F3Net Qian et al.
+(2020), NPR (Tan et al., 2024), CLIP (Radford et al., 2021), STIL (Gu et al., 2021a), XCLIP (Ni et al., 2022), VideoMAE (Tong et al., 2022), MINTIME (Coccomini et al., 2024), FTCN (Zheng et al., 2021), TALL (Xu et al., 2023a).
 Training settings.
@@
 Moreover, our DeMamba model can be effectively integrated into existing models, achieving significant improvements.
-For instance, by incorporating the DeMamba module into the XCLIP, the DeMamba-XCLIP-FT attains an average accuracy of 94.42% and an average AP of 97.10%, which represents a 10.22% increase in accuracy and 12.02% increase in AP compared to the original XCLIP.
+For example, integrating the DeMamba module into XCLIP results in DeMamba-XCLIP-FT achieving an average recall/F1/AP of 0.9392/0.9020/0.9710 respectively, which marks an improvement of 11.26%/17.72%/12.02% in recall/F1/AP over the original XCLIP.
 Note that PT (Partially Tuning) indicates that the backbone is frozen, with only the other parts being fine-tuned, while FT (Full Training) tuning the entire model.
@@
 Influence of scanning orders. As shown in Table 8, the continuous scan proposed in this paper effectively enhances performance compared to the traditional scanning method.
+Figure 3: Performance of training on scaled-up datasets on the testing set.
+Necessity of large volume of GenVideo for generalizability.
+To clarify the impact of data volume on detection performance, we conducted experiments using DeMamba on various scaled subsets of the GenVideo dataset.
+We designed two experiments to explore: 1) variation in the number of training samples within each generator: For the data generated by each generator, we randomly selected between 100 and 20,000 video samples to train the model.
+2) variation in the number of training samples across different generators: We randomly selected subsets to simulate scenarios with 1 to 10 generators in the dataset.
+In each scenario, we collected 20,000 video samples for the videos generated by each generator.
+Note that, for real videos, we randomly selected an equal number of samples to the total number of training generated video data to participate in the training.
+The test sets used in these experiments are the same as those mentioned in the main paper, they consist of the full test set videos, and the model selected for the experiment is DeMamba-XCLIP-FT.
+The comparative results of these experiments are detailed in Figure 4.
+The results clearly indicate a significant association between the scale of the dataset and the improvement in detection performance, further affirming the critical role of massive data in enhancing detection capabilities.
+Table 9: Training/inference time and QPS for different models.
+Model Training time (h) Inference time (s) QPS
+F3Net 40 188 2.47
+NPR 25.1 149 2.66
+STIL 11.7 174 2.42
+VideoMAE 52.5 283 1.32
+MINTIME 18.1 148 2.61
+FTCN 11.8 148 2.68
+TALL 6 150 2.71
+CLIP-B-FT 10.1 148 2.56
+DeMamba-CLIP-B-FT 14 150 2.54
+XCLIP-FT 11.5 149 2.55
+DeMamba-XCLIP-FT 15.1 153 2.53
+Training and inference efficiency.
+As detailed in Table 9, we report the training and inference times on 8 Tesla A100-80G GPUs, with QPS measured on a single GPU at batch size 1.
+Incorporating DeMamba into XCLIP does not significantly impact inference time compared to CLIP and XCLIP.
+Slower training is due to Mamba’s training characteristics, but it offers efficient inference post-training.
+Our method maintains competitive computational times relative to other video models, outperforming F3Net, STIL, and TALL.
 6 Broader impacts
@@
 Table 10: Model details and its performance on many-to-many generalization task.
-Model Detection level Input size Param (M) FLOPs(G) ACC AP
-CLIP-B Image $8\times 224\times 224$ 151.46 149.34 91.19 91.52
-NPR Image $8\times 224\times 224$ 1.44 14.08 85.30 82.45
-F3Net Image $8\times 299\times 299$ 48.31 145.04 83.45 88.73
-VideoMAE-B Video $16\times 224\times 224$ 86.54 147.52 88.61 94.54
-STIL Video $8\times 224\times 224$ 22.69 38.06 83.71 87.12
-XCLIP-B Video $8\times 224\times 224$ 121.26 141.28 85.66 86.77
-DeMamba-XCLIP Video $8\times 224\times 224$ 125.37 147.61 94.42 97.10
+Model Detection level Input size Param (M) FLOPs(G) R F1 AP
+CLIP-B Image $8\times 224\times 224$ 151.46 149.34 0.9057 0.7631 0.9152
+NPR Image $8\times 224\times 224$ 1.44 14.08 0.8408 0.7164 0.8245
+F3Net Image $8\times 299\times 299$ 48.31 145.04 0.8188 0.8024 0.8873
+VideoMAE-B Video $16\times 224\times 224$ 86.54 147.52 0.8750 0.8955 0.9454
+STIL Video $8\times 224\times 224$ 22.69 38.06 0.8222 0.7823 0.8712
+MINTIME-CLIP-B Video $8\times 224\times 224$ 201.35 221.0 0.8762 0.8231 0.9155
+FTCN-CLIP-B Video $8\times 224\times 224$ 174.82 146.69 0.8618 0.8755 0.9221
+TALL Video $8\times 224\times 224$ 86.92 15.2 0.8852 0.7419 0.8791
+XCLIP-B Video $8\times 224\times 224$ 121.26 141.28 0.6615 0.6497 0.7154
+DeMamba-XCLIP Video $8\times 224\times 224$ 125.37 147.61 0.9392 0.9020 0.9710
 Appendix B Experiment details